### v1

# Binarization-Aware Adjuster: Bridging Continuous Optimization and Binary Inference in Edge Detection

Abstract
Image edge detection (ED) faces a fundamental mismatch between training and inference: models are trained using continuous-valued outputs but evaluated using binary predictions. This misalignment, caused by the non-differentiability of binarization, weakens the link between learning objectives and actual task performance. In this paper, we propose a theoretical method to design a Binarization-Aware Adjuster (BAA), which explicitly incorporates binarization behavior into gradient-based optimization. At the core of BAA is a novel loss adjustment mechanism based on a Distance Weight Function (DWF), which reweights pixel-wise contributions according to their correctness and proximity to the decision boundary. This emphasizes decision-critical regions while down-weighting less influential ones. We also introduce a self-adaptive procedure to estimate the optimal binarization threshold for BAA, further aligning training dynamics with inference behavior. Extensive experiments across various architectures and datasets demonstrate the effectiveness of our approach. Beyond ED, BAA offers a generalizable strategy for bridging the gap between continuous optimization and discrete evaluation in structured prediction tasks.
I Introduction
Image edge detection (ED) is a fundamental task in computer vision, serving as a critical component for higher-level applications such as segmentation, inpainting, and object recognition [1, 2, 3]. While ED has advanced significantly—from early classical methods [4, 5] to modern deep learning-based approaches [6, 7, 8, 9, 10], a core theoretical inconsistency remains underexplored.
This inconsistency lies in the gap between the continuous-valued outputs used for training and the binary predictions required at inference. Ground-truth edge maps are binary, indicating the presence or absence of edges at each pixel. However, to enable gradient-based optimization, models are trained to produce continuous values in $[0,1]$, typically using loss functions such as binary cross-entropy (BCE). Because binarization is discontinuity, it is excluded from the training loop and applied only post hoc during evaluation. This disconnect can lead to suboptimal performance: a model that minimizes the training loss may not yield the most accurate binary predictions after thresholding.
In practice, a fixed threshold is used to convert predicted probabilities into a binary edge map. Predictions near this threshold are inherently uncertain and more prone to misclassification, making them especially critical to the final binary outcome. In contrast, confidently correct predictions, namely those far from the threshold and on the correct side, have limited impact on the binary result, but they often dominate the loss due to their ease of optimization. As a result, standard continuous-value loss functions allocate excessive learning capacity to uninformative regions while neglecting the more decision-critical predictions near the decision threshold. This imbalance impairs the effectiveness of training and may ultimately degrade binary performance.
While prior work has improved ED performance through architectural innovations [6, 7, 10] and alternative loss functions [11, 12, 13], few have addressed the theoretical misalignment between continuous-valued training objectives and binary evaluation metrics. The binarization step remains disconnected from the optimization process, leaving a fundamental gap in the learning framework.
To address this, we propose a theoretical method to design a binarization-aware adjuster (BAA) that explicitly incorporates binarization behavior into the training process. Central to our method is a novel threshold-aware loss adjustment mechanism that dynamically modulates each pixel’s contribution to the loss based on its proximity to the decision threshold and correctness. Predictions near the threshold or those misclassified receive higher weights, while confidently correct predictions are down-weighted. This principled adjustment redirects the learning signal toward decision-critical regions, aligning optimization objectives with the binary nature of the task.
Our main contributions are as follows:
  • We identify a core mismatch in ED: the disconnect between continuous-valued supervision in model training and binary-valued final output in practical usage, and propose the BAA method to resolve it.
  • We introduce a principled method for constructing BAA based on Distance Weight Function (DWF), and demonstrate how it generalizes and improves upon loss formulations.
  • We validate our approach through extensive experiments across multiple benchmark datasets and architectures, demonstrating the effectiveness of the proposed method.
II Previous Works
This section reviews the landscape of ED research across datasets, model architectures, loss functions, and evaluation protocols, with an emphasis on foundational developments and persisting challenges.
II-A Datasets
Early ED research relied heavily on datasets designed for related tasks such as contour detection and image segmentation. Notable benchmarks like BSDS300 and BSDS500 [14], featuring multi-annotator contour labels, remain widely used. Other commonly adopted datasets include MDBD [15] for boundary detection, NYUD [16] for semantic segmentation, and general-purpose datasets such as PASCAL VOC [17], COCO [18], SceneParse150 [19], and Cityscapes [20]. However, these were not originally designed for ED.
As the field matured, specialized ED datasets emerged. BIPED and its improved version BIPED2 [21] introduced high-resolution single-edge annotations. BRIND [6] extended the annotations of BSDS500 with semantically labeled edges, including reflectance, illumination, normal, and depth. UDED [22] offered fewer samples but prioritized annotation quality and consistency, enabling more reliable evaluation.
Nevertheless, annotation quality remains a key bottleneck. Human-labeled edge maps often suffer from noise, inconsistency, and semantic ambiguity. Although several studies have proposed more robust training strategies and refined labeling protocols [23, 24, 25], annotation noise continues to challenge both model generalization and fair benchmarking.
II-B Model Architectures
The evolution of ED models can be divided into three phases: classical heuristics, early learning-based methods, and deep learning-driven approaches.
Classical methods such as the Sobel operator [4] and Canny detector [5] used hand-crafted gradient filters to capture intensity discontinuities. While computationally efficient, these methods lack adaptability to semantic structures and struggle in complex visual environments.
Early learning-based models incorporated statistical learning using manually crafted features like texture descriptors [26], sketch tokens [27], or histogram-based cues [28], paired with classifiers such as structured forests [29] or logistic regression [30]. However, these approaches were fundamentally limited by their reliance on hand-engineered features and had limited generalization across diverse scenes.
The advent of deep learning, particularly convolutional neural networks (CNNs), opened a new path for ED. HED [11] introduced deep supervision across network layers from multiple resolutions, while RCF [31] and BDCN [32] refined the multi-scale fusion techniques. Models such as PiDiNet [33] and DexiNet [34] focused on designing new structures for improving ED prediction. Beyond CNNs, transformer-based methods [7, 35] and diffusion-based approaches [8, 9] explore global context modeling and have also been employed in ED-related tasks, though their advantages in the ED task itself remain under study. Recently, the Extractor-Selector (E-S) paradigm [10] introduced a two-stage framework combining feature extraction with a learnable selection module, offering a new perspective on the challenge of discriminative feature selection.
II-C Loss Functions
Loss function design plays a critical role in ED. The Weighted Binary Cross-Entropy (WBCE) loss remains the most widely adopted approach, emphasizing edge pixels to address the low recall rate of the BCE loss. Despite its effectiveness in obtaining a high recall rate, WBCE often requires non-maximum suppression (NMS) at inference time to obtain clean, thin edge maps, resulting in a trade-off between quantitative score and prediction quality.
To address this, various alternative loss functions have been proposed. Dice loss [36] emphasizes region-level overlap, while tracing loss [37] targets the proportion of edges within edge-adjacent areas. Ranking-based approaches, such as RankED [12] and AP-loss [38], aim to enhance precision by optimizing pixel-level ranking. More recently, SWBCE [13] accounts for human uncertainty by rebalancing the contributions of edge and non-edge regions.
While these advanced losses provide perceptual or structural improvements, they often introduce new challenges such as instability, reduced precision, or sensitivity to hyperparameters, hindering their widespread use. A unified loss that aligns training supervision with binary evaluation, while maintaining stability and generalization, remains an open problem.
II-D Evaluation Metrics
Standard evaluation metrics for ED were established by [26] and remain widely adopted. The two most common metrics are:
  • ODS (Optimal Dataset Score): Best F1-score computed using a single optimal binarization threshold applied across the entire dataset.
  • OIS (Optimal Image Score): Mean of best F1-scores computed per image using image-specific optimal thresholds.
These metrics are computed using an error-tolerance matching strategy to allow minor spatial misalignments. However, the error-tolerance radius is often dataset-specific and can influence results and comparability. Early works did not unify the error-tolerance benchmark among datasets until recent studies advocated for a fixed 1-pixel tolerance to ensure stricter, more consistent evaluation [25].
III Methodology
III-A Motivation and Insight
A key challenge in training ED models is the mismatch between training objectives and evaluation protocols. During training, ED models generate continuous-valued outputs in the range $[0,1]$, typically optimized using pixel-wise loss functions such as WBCE. In contrast, evaluation metrics are computed on binarized edge maps, obtained by applying a fixed threshold to these outputs. Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process. This disconnect creates a fundamental issue: the loss function does not explicitly align with the discrete classification decisions that ultimately determine evaluation performance.
This gap introduces a critical inefficiency in optimization. Pixels whose predicted values lie on the correct side and far from the decision threshold, namely those correct pixels with high confidence, tend to dominate the gradient signal since they are easy to optimize. In contrast, pixels with predictions near the threshold, which are more ambiguous and critical to classification, contribute less to the gradients since they are more difficult to optimize. Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority. This misallocation of optimization effort can prevent the optimal parameter be learned, resulting in suboptimal learning and degraded final performance.
For example, Table I reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8. Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss. Moreover, such confident correct predictions are often easy to optimize. Indeed, it is because they are easy to optimize that they become the correctly predicted ones with high confidence. Therefore, training naturally prioritizes them for their optimization ease, even though further improvement on them yields no benefit for the final decision. In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold. Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
TABLE I: Optimal binarization thresholds under WBCE training: Evaluation uses 1-pixel error tolerance without NMS. Experimental settings follow the ones in Section IV-A. Models are retrained on the pre-training set and evaluated on the validation set.
UDED BSDS500 BRIND BIPED2
HED 0.74 0.85 0.77 0.79
BDCN 0.65 0.79 0.78 0.72
Dexi 0.63 0.87 0.79 0.79
EES3-Double 0.61 0.86 0.77 0.80
To bridge the continuous optimization and the binarization inference, we propose a novel loss adjustment mechanism, realized by a BAA, that dynamically re-weights pixel-wise loss contributions based on both proximity to the decision threshold and classification correctness. By explicitly incorporating binarization behavior into the training process, The BAA method ensures that the model allocates its learning capacity to the most decision-critical pixels.
III-B Adjusted Loss Function Formulation
Although designing the BAA is non-trivial, the idea of using an adjuster is conceptually simple. Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
We define an adjusted loss as a loss that incorporates a pixel-wise weighting factor:
$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(P_{i},G_{i})+\delta)L(P_{i},G_{i})$ (1)
where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single pixel ($L(P_{i},G_{i})$) and on the whole image ($L(Pred,Gt)$)., $Pred$ and $Gt$ represent the model prediction and the ground truth, respectively, $P_{i}$ and $G_{i}$ represent the model prediction and the ground truth of the $i$-th pixel, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the adjusted loss to stabilize training. The function $Adj$ is called an Adjuster, where $Adj(P_{i},G_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(P_{i},G_{i})$ is the binary mask $\hat{Adj}_{thr}(P_{i},G_{i})$, defined as:
$\hat{Adj}_{thr}(P_{i},G_{i})=\left\{\begin{matrix}&0&If\ (P_{i}-thr)(G_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right.$ (2)
where $thr$ is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
  • If a pixel is correctly classified, then the loss on it is 0.
  • If a pixel is incorrectly classified, then the loss on it is positive.
This adjuster emphasizes only misclassified pixels. However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization. Therefore, a smoother, differentiable formulation is necessary.
III-C Design of the Binarization-Aware Adjuster
Let $thr$ denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
  • $C1\ (Continuity)$: The adjuster $Adj$ must be continuous to allow gradient-based optimization.
  • $C2\ (Non-negativity)$, $Adj(P_{i},G_{i})$ should be non-negative since it represents the weight of a pixel. Without loss generality, assume that $Adj(P_{i},G_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
  • $C3\ (Uncertainty\ Emphasis)$: If a pixel is correctly classified, then the closer the prediction is to $thr$, the higher the weight $Adj(P_{i},G_{i})$ of it should be, reflecting the loss caused by the uncertainty.
  • $C4\ (Confidence\ Saturation)$: If a pixel is correctly classified with high confidence (i.e., prediction far from $thr$), then $Adj(P_{i},G_{i})$ should decay to 0 quickly as $P_{i}$ goes away from $thr$, representing that further optimizing provide little benefits. Indeed, $Adj(P_{i},G_{i})$ should close to the maximal value when $P_{i}$ close to $thr$, remain large when $P_{i}$ is not far from $thr$, and decay to 0 after some pre-defined distance $thr\_dev$.
  • $C5\ (Classification\ Sensitivity)$: If a pixel is misclassified, namely if it is an edge one but $P_{i}<thr$ or is a non-edge one but $P_{i}\geq thr$, its weight $Adj(P_{i},G_{i})$ should be maximal, representing correctly classification is of the first priority.
These conditions ensure that misclassified pixels always receive full attention, and optimization of correct pixels is focused on the uncertain parts while conserving effort on highly confident correct parts. We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short. Namely:
Definition 1
A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and $thr\_dev>0$ is a function $[0,1]^{2}\longrightarrow[0,1]$ satisfying the above conditions C1 to C5.
Let $thr\_dev$ denote the threshold window from $C4$. To construct a BAA, we define that:
Definition 2
A function $f_{thr\_dev}(x)$ is said to be a Distance Weight Function (DWF) regarding the threshold window $thr\_dev\in\mathbb{R}_{+}$, if it satisfies the following properties:
  • Continuity: $f_{thr\_dev}(x):[0,thr\_dev]\longrightarrow[0,1]$ is continuous.
  • Monotonically Decreasing: $f_{thr\_dev}(x_{1})\leq f_{thr\_dev}(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ with $x_{1}\geq x_{2}$. (or equivalently, $\frac{df_{thr\_dev}}{dx}(x)\leq 0$ for differentiable $f_{the\_dev}(x)$).
  • Concavity: $f(\lambda x_{1}+(1-\lambda)x_{2})\geq\lambda f(x_{1})+(1-\lambda)f(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ and $\lambda\in[0,1]$. (or equivalently, $\frac{d^{2}f_{thr\_dev}}{dx^{2}}(x)\leq 0$ for secondary differentiable $f_{the\_dev}(x)$).
  • Boundary Conditions: $f_{thr\_dev}(0)=1$, and $f_{thr\_dev}(thr\_dev)=0$.
Then:
$\bar{Adj}(P_{i},G_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|P_{i}-thr|)&0\leq|P_{i}-thr|\leq thr\_dev\\ &0\ &|P_{i}-thr|>thr\_dev\end{matrix}\right.$ (3)
satisfies $C1-C4$.
An example of the DWF $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ can be constructed as:
$f_{b,thr\_dev}(x)=\frac{e^{b\times x}-e^{b\times thr\_dev}}{1-e^{b\times thr\_dev}}$ (4)
where $b\geq 0$ is pre-defined, controlling the decay rate.
To also meet $C5$, we define the masked distance as:
Definition 3
The ($[0,1]$) masked distance $MD_{thr}(x,y):[0,1]^{2}\longrightarrow[0,1]$ regarding the threshold $thr\in[0,1]$ is defined as:
$MD_{thr}(x,y)=(x-thr)\times y+(thr-x)\times(1-y)$ (5)
Hence, $MD_{thr}(P_{i},G_{i})$ measures both the correctness and confidence of the pixel $i$. For a wrongly predicted pixel $i$, namely whose prediction is on the different side of the $thr$ as its ground truth value, $MD_{thr}(P_{i},G_{i})\leq 0$, while for a correctly predicted pixel $i$, $MD_{thr}(P_{i},G_{i})$ is the $L^{1}$ distance of $P_{i}$ from the binarization threshold $thr$.
Now, a BAA can be defined as:
$BAA_{thr,thr\_dev}(P_{i},G_{i})=\hat{f}_{thr\_dev}(MD_{thr}(P_{i},G_{i}))$ (6)
where $\hat{f}_{thr\_dev}$ is an extension of a selected DWF $f_{thr\_dev}(x)$, defined as:
$\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<0\\ &f_{thr\_dev}(x)&0\leq x\leq thr\_dev\\ &0&x>thr\_dev\end{matrix}\right.$ (7)
III-D Theoretical Properties of the BAA
To better understand the proposed adjuster’s behavior, we analyze its mathematical properties. Let $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ be defined as in Equations 4 and $\hat{f}_{thr\_dev}(x)=\hat{f}_{b,thr\_dev}(x)$ be its extension defined in Equation 7. Then $\hat{f}_{b,thr\_dev}(x)$ serves as a smooth approximation to a hard binarization function. Specifically, it generalizes the discrete threshold-based binary weighting function $B_{thr\_dev}(x)$ on threshold $thr\_dev$, defined as:
$B_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<thr\_dev\\ &0\ &x\geq thr\_dev\end{matrix}\right.$ (8)
The two functions coincide as $b\rightarrow+\infty$:
$\lim_{b\rightarrow+\infty}\hat{f}_{b,thr\_dev}(x)=B_{thr\_dev}(x)$ (9)
Therefore, the extended function $\hat{f}_{b,thr\_dev}$ is a continuous approximation for threshold logic $B_{the\_dev}(x)$. Accordingly, the map $BAA_{thr,thr\_dev}(P_{i},G_{i})=BAA_{thr,b,thr\_dev}(P_{i},G_{i})$, defined by Equation 6, converges to a discrete decision-based adjuster $\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$, defined as follow: {strip}
$\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})=\left\{\begin{matrix}&0&(P_{i}-thr)(G_{i}-thr)\geq 0\ \&\ |P_{i}-thr|\geq thr\_dev\\ &1&Otherwise\end{matrix}\right.$ (10)
i.e.,
$\small\lim_{b\rightarrow+\infty}BAA_{thr,b,thr\_dev}(P_{i},G_{i})=\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$ (11)
This confirms that the BAA bridges the binarization logic with the continuous-loss gradient-based training. Furthermore, let also $thr\_dev\rightarrow 0^{+}$, then it is reduced to $\hat{Adj}_{thr}$ from Equation 2.
Moreover, for a DWF $f_{thr\_dev}(x)$, the following bounds holds:
$1-\frac{x}{thr\_dev}\leq f_{thr\_dev}(x)\leq 1$ (12)
Therefore:
$\lim_{thr\_dev\rightarrow+\infty}f_{thr\_dev}(x)=1$ (13)
which implies that:
$\lim_{thr\_dev\rightarrow+\infty}BAA_{thr,thr\_dev}(x)=1$ (14)
and the adjusted loss $L_{Adj,\delta}$ defined in Equation 1 degenerates to the baseline loss $L$ up to a constant:
$\lim_{thr\_dev\to\infty}L_{Adj,\delta}(Pred,Gt)=(1+\delta)L(Pred,Gt)$ (15)
In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 4, generalizes the baseline loss $L$. It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided.. These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation in ED.
III-E Hyperparameter Selection and Training Protocol
The proposed $BAA$ depends on three hyperparameters: $thr$: binarization threshold, $thr\_dev$: threshold window beyond which correct predictions are ignored in optimization, and $b$ (when employing $f_{b,thr\_dev}(x)$ from Equation 4 as the DWF): DWF decay rate,
The DWF decay rate $b$ occurs specifically in the DWF defined in Equation 4, controlling the concavity of $f_{b,thr\_dev}(x)$, determining how sharply the adjusted weight decays as predictions diverge from the decision boundary $thr$. Its effect is illustrated in Figure 1, and we set $b=16$ as the default.
Figure 1:  The different choices of $b$ in the DWF $f_{b,thr\_dev}(x)$ presented in Equation 4
$thr\_dev$ defines the margin around $thr$ within which a correct prediction is considered uncertain. A correct prediction with $|P_{i}-thr|>thr\_dev$ is considered as confidently correct and receives zero weight. We empirically choose $thr\_dev=0.2$.
$thr$ is the most critical hyperparameter in the BAA. Experiments (Table I) show that the optimal value for ED models typically lies between 0.6 and 0.8, so we use $thr=0.7$ as the default.
However, given its importance, it is worth providing a calibration method for it. Specifically, we suggest a self-adaptive $thr$ chosen by pre-training: pre-train the model on a sub-dataset without applying BAA and evaluate the pre-trained model on a validation set to select a suitable $thr$ for BAA.
Therefore, the final training protocol of an ED model $M$ under a selected baseline loss function $L$ on a dataset $D$, with the presented BAA method can be formulated as:
  • Step 1: Split $D$ into pre-train set $PT$, validation set $V$, and test set $T$.
  • Step 2: Train $M$ on $PT$ using baseline loss $L$, and evaluate on $V$ to estimate $thr$.
  • Step 3: Retrain $M$ via $L_{BAA,\delta}$, on $PT\cup V$, and evaluate finally on $T$.
In experiments, $\delta$ is simply set to be 1.
IV Experiment
This section presents experiments designed to evaluate the proposed method.
IV-A Experimental Setting
To assess the effectiveness of the proposed binarization adjustment mechanism, we adopt the adjuster designed by Equation 4, 5, 6, and 7 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
IV-A1 Baseline and Benchmark
The proposed method is evaluated across a range of ED models, including:
  • HED (2015) [11].
  • BDCN (2022) [32].
  • Dexi (2023) [34].
  • EES3-Double (2025) [10], the EES framework, integrating the above three models with the double-size selector.
Experiments are conducted on four benchmark datasets:
  • BRIND [6].
  • BSDS500 [14].
  • UDED [22].
  • BIPED2 [21].
All models are retrained from scratch on each dataset to ensure fairness. For consistency and simplicity, only the final output of each model is used for supervision, bypassing any need for hyperparameter tuning in intermediate layers. Each model is trained under the following three configurations of loss function for comparison:
  • Standard WBCE (baseline).
  • WBCE with a fixed-threshold BAA using $thr=0.7$ (denoted as BAA-0.7)
  • WBCE with a self-adaptive adjuster where $thr$ is automatically selected following the method in Section III-E (denoted as BAA-SA).
Unless otherwise specified, the hyperparameters for the adjusted loss defined by the BAA are fixed at $thr\_dev=0.2$, $b=16$, and $\delta=1$.
Evaluation protocols follow the standard ED metrics (ODS and OIS) introduced in [26], under the strictest 1-pixel error tolerance without post-processing techniques such as non-maximal-suppression (NMS), suggested in [25]. The evaluations under the traditional relaxed error tolerance (4.3 to 11.1 pixels) with NMS are provided in the appendix.
IV-A2 Dataset Partitions and Augmentation
Dataset preprocessing and postprocessing procedures follow those introduced in [10]. Details are stated as follows.
The dataset partitions are:
  • BRIND and BSDS500: 400 training images, 100 testing images.
  • UDED: 20 training images, 7 testing images (excluding two images with resolution below $320\times 320$).
  • BIPED2: 200 training images, 50 testing images.
For training under the self-adaptive adjuster setting in section III-E, each training set is further divided into pre-training and validation subsets:
  • BRIND and BSDS500: 300 pre-training, 100 validation images.
  • UDED: 15 pre-training, 5 validation images.
  • BIPED2: 150 pre-training, 50 validation images.
Data augmentations are applied, including:
  • Halving the images until both height and width are strictly below 640 pixels.
  • Rotations at four angles, $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, with horizontal flips.
  • Addition of noiseless data.
IV-A3 Training Note
During training, input images are randomly cropped to $320\times 320$ and refreshed every 5 epochs. The batch size is set to 8. All models are trained on an A100 GPU using Adam optimizer, with a learning rate of 0.0001 and a weight decay of $10^{-8}$.
Training runs for 200 epochs on UDED and 50 epochs on the remaining datasets.
IV-A4 Inference Method
In the inference phase, each image is split into overlapping $320\times 320$ patches with a stride of 304 pixels (i.e., 16 pixels overlap). The model predicts edge maps on each patch independently, and the outputs are aggregated to form the final edge map.
Further implementation details can be found in the codes or in [10].
IV-B Experimental Result
Tables II and III report the ED performance of all methods on BRIND, BSDS500, UDED, and BIPED2 datasets under the strict 1-pixel error tolerance setting, without applying NMS. The evaluation includes ODS and OIS scores for each configuration.
TABLE II: Results on BRIND and BSDS500 under 1-pixel error tolerance without NMS: For BRIND, all annotations per image are merged into a single binary ground truth map, where a pixel is considered an edge one if it is marked as such by any annotator. For BSDS500, only label 1 is used as the ground truth. Notations follow those introduced in the main-text. The best scores for each model are highlighted in bold. Rows labeled A, A-BAA-0.7, and A-BAA-SA represent the average performance across the four models, while Improvement rows indicate the relative improvements over the baseline.
BRIND BSDS500
ODS OIS ODS OIS
HED (2015) 0.664 0.673 0.444 0.455
HED-BAA-0.7 0.668 0.676 0.441 0.453
HED-BAA-SA 0.664 0.672 0.455 0.465
BDCN (2022) 0.655 0.663 0.432 0.446
BDCN-BAA-0.7 0.662 0.671 0.445 0.452
BDCN-BAA-SA 0.665 0.674 0.454 0.468
Dexi (2023) 0.675 0.683 0.465 0.473
Dexi-BAA-0.7 0.674 0.683 0.475 0.485
Dexi-BAA-SA 0.679 0.687 0.470 0.480
EES3-Double (2025) 0.683 0.690 0.488 0.504
EES3-Double-BAA-0.7 0.690 0.697 0.493 0.509
EES3-Double-BAA-SA 0.690 0.697 0.497 0.507
A 0.669 0.677 0.457 0.470
A-BAA-0.7 0.674 0.682 0.464 0.475
[…]
A-BAA-SA 0.675 0.683 0.469 0.480
[…]
TABLE III: Results on UDED and BIPED2 under 1-pixel error tolerance without NMS: Notations follow the ones in TableII.
UDED BIPED2
ODS OIS ODS OIS
HED (2015) 0.709 0.742 0.630 0.633
HED-BAA-0.7 0.727 0.755 0.629 0.634
HED-BAA-SA 0.712 0.748 0.633 0.635
BDCN (2022) 0.688 0.723 0.636 0.639
BDCN-BAA-0.7 0.713 0.741 0.637 0.639
BDCN-BAA-SA 0.712 0.749 0.640 0.643
Dexi (2023) 0.700 0.738 0.643 0.648
Dexi-BAA-0.7 0.721 0.752 0.640 0.646
Dexi-BAA-SA 0.731 0.765 0.641 0.646
EES3-Double (2025) 0.751 0.780 0.670 0.673
EES3-Double-BAA-0.7 0.762 0.794 0.671 0.675
EES3-Double-BAA-SA 0.758 0.791 0.675 0.680
A 0.712 0.746 0.645 0.648
A-BAA-0.7 0.731 0.761 0.644 0.649
[…]
A-BAA-SA 0.729 0.763 0.647 0.651
[…]
The results clearly demonstrate that incorporating the proposed BAA method, whether with the fixed threshold (BAA-0.7) or the self-adaptive strategy (BAA-SA), leads to improvements in performance across most models and datasets.
Both variants of the BAA loss improve the baseline WBCE loss across all models and datasets. The self-adaptive version (BAA-SA) tends to achieve slightly better performance than the fixed threshold in most cases.
[…]
These results strongly validate the effectiveness of the proposed BAA method, especially the self-adaptive variant. The method is robust across different backbone architectures and datasets, offering consistent improvements with minimal modifications to the existing training pipeline.
IV-C Ablation Study
Figure 2 illustrates how ODS and OIS scores vary with different values of $thr\_dev$ (from 0.1 to 0.7) and decay rates $b$ (8, 16, 32) in the DWF (Equation 4). All experiments are conducted on the BRIND dataset using the EES3-Double model trained with the BAA-WBCE loss. The threshold $thr$ is fixed at 0.7, and $\delta$ at 1.
This ablation study shows that across all tested configurations, $thr\_dev=0.2$ consistently yields the best performance. Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
Figure 2:  ODS and OIS across $thr\_dev$: Figures show how ODS and OIS vary with $thr\_dev$ values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32). Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss. Threshold $thr$ is fixed at 0.7 and $\delta$ at 1.
V Conclusion
This paper tackles a fundamental yet underexplored issue in image edge detection (ED): the misalignment between the continuous-valued outputs optimized during training and the binary predictions required at inference and evaluation. This discrepancy, often overlooked due to the discontinuity of binarization, undermines the alignment between training objectives and final task performance.
To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process. Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each pixel’s contribution to the loss based on its prediction correctness and confidence. It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision. The resulting adjusted loss function from the BAA generalizes traditional loss formulations and embeds binarization behavior directly into the optimization process. In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior.
Extensive experiments across diverse models and datasets validate the effectiveness and robustness of the proposed method. Beyond ED, our BAA approach can also provide a theoretical method for bridging the gap between continuous optimization and discrete evaluation, offering a promising direction for broader applications in structured prediction tasks such as segmentation and classification.

### v2

@@
-Binarization-Aware Adjuster: Bridging Continuous Optimization and Binary Inference in Edge Detection
+Binarization-Aware Adjuster: A Theoretical Framework for Bridging Continuous Optimization and Discrete Inference with Application to Edge Detection
 Abstract
-Image edge detection (ED) faces a fundamental mismatch between training and inference: models are trained using continuous-valued outputs but evaluated using binary predictions.
-This misalignment, caused by the non-differentiability of binarization, weakens the link between learning objectives and actual task performance.
-In this paper, we propose a theoretical method to design a Binarization-Aware Adjuster (BAA), which explicitly incorporates binarization behavior into gradient-based optimization.
-At the core of BAA is a novel loss adjustment mechanism based on a Distance Weight Function (DWF), which reweights pixel-wise contributions according to their correctness and proximity to the decision boundary.
-This emphasizes decision-critical regions while down-weighting less influential ones.
-We also introduce a self-adaptive procedure to estimate the optimal binarization threshold for BAA, further aligning training dynamics with inference behavior.
-Extensive experiments across various architectures and datasets demonstrate the effectiveness of our approach.
-Beyond ED, BAA offers a generalizable strategy for bridging the gap between continuous optimization and discrete evaluation in structured prediction tasks.
-I Introduction
-Image edge detection (ED) is a fundamental task in computer vision, serving as a critical component for higher-level applications such as segmentation, inpainting, and object recognition [1, 2, 3].
-While ED has advanced significantly—from early classical methods [4, 5] to modern deep learning-based approaches [6, 7, 8, 9, 10], a core theoretical inconsistency remains underexplored.
-This inconsistency lies in the gap between the continuous-valued outputs used for training and the binary predictions required at inference.
-Ground-truth edge maps are binary, indicating the presence or absence of edges at each pixel.
-However, to enable gradient-based optimization, models are trained to produce continuous values in $[0,1]$, typically using loss functions such as binary cross-entropy (BCE).
-Because binarization is discontinuity, it is excluded from the training loop and applied only post hoc during evaluation.
-This disconnect can lead to suboptimal performance: a model that minimizes the training loss may not yield the most accurate binary predictions after thresholding.
-In practice, a fixed threshold is used to convert predicted probabilities into a binary edge map.
-Predictions near this threshold are inherently uncertain and more prone to misclassification, making them especially critical to the final binary outcome.
-In contrast, confidently correct predictions, namely those far from the threshold and on the correct side, have limited impact on the binary result, but they often dominate the loss due to their ease of optimization.
-As a result, standard continuous-value loss functions allocate excessive learning capacity to uninformative regions while neglecting the more decision-critical predictions near the decision threshold.
-This imbalance impairs the effectiveness of training and may ultimately degrade binary performance.
-While prior work has improved ED performance through architectural innovations [6, 7, 10] and alternative loss functions [11, 12, 13], few have addressed the theoretical misalignment between continuous-valued training objectives and binary evaluation metrics.
-The binarization step remains disconnected from the optimization process, leaving a fundamental gap in the learning framework.
-To address this, we propose a theoretical method to design a binarization-aware adjuster (BAA) that explicitly incorporates binarization behavior into the training process.
-Central to our method is a novel threshold-aware loss adjustment mechanism that dynamically modulates each pixel’s contribution to the loss based on its proximity to the decision threshold and correctness.
-Predictions near the threshold or those misclassified receive higher weights, while confidently correct predictions are down-weighted.
-This principled adjustment redirects the learning signal toward decision-critical regions, aligning optimization objectives with the binary nature of the task.
-Our main contributions are as follows:
-  • We identify a core mismatch in ED: the disconnect between continuous-valued supervision in model training and binary-valued final output in practical usage, and propose the BAA method to resolve it.
-  • We introduce a principled method for constructing BAA based on Distance Weight Function (DWF), and demonstrate how it generalizes and improves upon loss formulations.
-  • We validate our approach through extensive experiments across multiple benchmark datasets and architectures, demonstrating the effectiveness of the proposed method.
-II Previous Works
-This section reviews the landscape of ED research across datasets, model architectures, loss functions, and evaluation protocols, with an emphasis on foundational developments and persisting challenges.
-II-A Datasets
-Early ED research relied heavily on datasets designed for related tasks such as contour detection and image segmentation.
-Notable benchmarks like BSDS300 and BSDS500 [14], featuring multi-annotator contour labels, remain widely used.
-Other commonly adopted datasets include MDBD [15] for boundary detection, NYUD [16] for semantic segmentation, and general-purpose datasets such as PASCAL VOC [17], COCO [18], SceneParse150 [19], and Cityscapes [20].
-However, these were not originally designed for ED.
-As the field matured, specialized ED datasets emerged.
-BIPED and its improved version BIPED2 [21] introduced high-resolution single-edge annotations.
-BRIND [6] extended the annotations of BSDS500 with semantically labeled edges, including reflectance, illumination, normal, and depth.
-UDED [22] offered fewer samples but prioritized annotation quality and consistency, enabling more reliable evaluation.
-Nevertheless, annotation quality remains a key bottleneck.
-Human-labeled edge maps often suffer from noise, inconsistency, and semantic ambiguity.
-Although several studies have proposed more robust training strategies and refined labeling protocols [23, 24, 25], annotation noise continues to challenge both model generalization and fair benchmarking.
-II-B Model Architectures
-The evolution of ED models can be divided into three phases: classical heuristics, early learning-based methods, and deep learning-driven approaches.
-Classical methods such as the Sobel operator [4] and Canny detector [5] used hand-crafted gradient filters to capture intensity discontinuities.
-While computationally efficient, these methods lack adaptability to semantic structures and struggle in complex visual environments.
-Early learning-based models incorporated statistical learning using manually crafted features like texture descriptors [26], sketch tokens [27], or histogram-based cues [28], paired with classifiers such as structured forests [29] or logistic regression [30].
-However, these approaches were fundamentally limited by their reliance on hand-engineered features and had limited generalization across diverse scenes.
-The advent of deep learning, particularly convolutional neural networks (CNNs), opened a new path for ED.
-HED [11] introduced deep supervision across network layers from multiple resolutions, while RCF [31] and BDCN [32] refined the multi-scale fusion techniques.
-Models such as PiDiNet [33] and DexiNet [34] focused on designing new structures for improving ED prediction.
-Beyond CNNs, transformer-based methods [7, 35] and diffusion-based approaches [8, 9] explore global context modeling and have also been employed in ED-related tasks, though their advantages in the ED task itself remain under study.
-Recently, the Extractor-Selector (E-S) paradigm [10] introduced a two-stage framework combining feature extraction with a learnable selection module, offering a new perspective on the challenge of discriminative feature selection.
-II-C Loss Functions
-Loss function design plays a critical role in ED.
-The Weighted Binary Cross-Entropy (WBCE) loss remains the most widely adopted approach, emphasizing edge pixels to address the low recall rate of the BCE loss.
-Despite its effectiveness in obtaining a high recall rate, WBCE often requires non-maximum suppression (NMS) at inference time to obtain clean, thin edge maps, resulting in a trade-off between quantitative score and prediction quality.
-To address this, various alternative loss functions have been proposed.
-Dice loss [36] emphasizes region-level overlap, while tracing loss [37] targets the proportion of edges within edge-adjacent areas.
-Ranking-based approaches, such as RankED [12] and AP-loss [38], aim to enhance precision by optimizing pixel-level ranking.
-More recently, SWBCE [13] accounts for human uncertainty by rebalancing the contributions of edge and non-edge regions.
-While these advanced losses provide perceptual or structural improvements, they often introduce new challenges such as instability, reduced precision, or sensitivity to hyperparameters, hindering their widespread use.
-A unified loss that aligns training supervision with binary evaluation, while maintaining stability and generalization, remains an open problem.
-II-D Evaluation Metrics
-Standard evaluation metrics for ED were established by [26] and remain widely adopted. The two most common metrics are:
-  • ODS (Optimal Dataset Score): Best F1-score computed using a single optimal binarization threshold applied across the entire dataset.
-  • OIS (Optimal Image Score): Mean of best F1-scores computed per image using image-specific optimal thresholds.
-These metrics are computed using an error-tolerance matching strategy to allow minor spatial misalignments.
-However, the error-tolerance radius is often dataset-specific and can influence results and comparability.
-Early works did not unify the error-tolerance benchmark among datasets until recent studies advocated for a fixed 1-pixel tolerance to ensure stricter, more consistent evaluation [25].
-III Methodology
-III-A Motivation and Insight
-A key challenge in training ED models is the mismatch between training objectives and evaluation protocols.
-During training, ED models generate continuous-valued outputs in the range $[0,1]$, typically optimized using pixel-wise loss functions such as WBCE.
-In contrast, evaluation metrics are computed on binarized edge maps, obtained by applying a fixed threshold to these outputs.
+In machine learning, discrete decision-making tasks exhibit a fundamental inconsistency between training and inference: models are optimized using continuous-valued outputs, yet evaluated through discrete predictions.
+This discrepancy arises from the non-differentiability of discretization operations, weakening the alignment between optimization objectives and practical decision outcomes.
+To address this, we present a theoretical framework for constructing a Binarization-Aware Adjuster (BAA) that integrates binarization behavior directly into gradient-based learning.
+Central to the approach is a Distance Weight Function (DWF) that dynamically modulates pixel-wise loss contributions based on prediction correctness and proximity to the decision boundary, thereby emphasizing decision-critical regions while de-emphasizing confidently correct samples.
+Furthermore, a self-adaptive threshold estimation procedure is introduced to better match optimization dynamics with inference conditions.
+As one of its applications, we implement experiments on the edge detection (ED) task, which also demonstrate the effectiveness of the proposed method experimentally.
+Beyond binary decision tasks and ED, the proposed framework provides a general strategy for aligning continuous optimization with discrete evaluation and can be extended to multi-valued decision processes in broader structured prediction problems.
+keywords:
+Adjusted Loss Function, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
+[…]
+[…]
+1 Introduction
+Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately require discrete decisions (e.g., binary or categorical outputs).
+However, learning methods, such as deep neural networks, constrained by the need for differentiable computation, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks [6, 7, 8, 9, 10].
+As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a fundamental gap arises between continuous-valued optimization during training and the discrete inference deployed at test time.
+This gap may lead to suboptimal performance, as optimization does not always align with decision-critical behavior.
+This mismatch is particularly evident in pixel-level binary prediction tasks.
+For example, edge detection (ED) aims to generate binary edge maps but is typically trained using continuous-valued loss functions such as binary cross-entropy (BCE).
+Since binarization is non-differentiable, it is excluded from the optimization loop and applied only post hoc during evaluation.
+As a result, model updates tend to prioritize pixels that are easy to optimize but have limited influence on the binary output, whereas pixels near the decision boundary, which are more critical for final accuracy, receive insufficient emphasis due to their lower weights and higher optimization difficulty.
+Several studies have attempted to mitigate the inconsistency between decision behavior and training objectives, typically by changing losses that rebalance decision classes, such as weighted BCE [11] or alternative ones [12, 13].
+However, to the best of our knowledge, few existing methods explicitly integrate binarization (or discretization) behavior into gradient-based optimization, leaving the fundamental optimization–inference misalignment unresolved.
+To address this issue, we propose a theoretically grounded framework named Binarization-Aware Adjuster (BAA), which explicitly embeds binarization behavior into the optimization process.
+The core idea is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which dynamically reweights pixel-level loss according to prediction correctness and proximity to the decision threshold.
+This enables the model to allocate greater learning capacity to decision-critical regions while down-weighting confidently correct predictions that exert limited influence on binary outputs.
+Additionally, a self-adaptive threshold estimation strategy is introduced to align the adjuster with inference behavior.
+Since the BAA is continuous and differentiable, it can be seamlessly incorporated into existing gradient-based optimization pipelines, allowing discrete inference behavior to influence learning directly.
+The proposed framework is designed for binary decision tasks and is applied to the ED task to demonstrate its effectiveness.
+ED is selected as a representative implementation scenario due to its foundational role in low-level vision and its strong dependence on precise decision boundaries.
+Furthermore, many structured prediction tasks can be viewed as extensions or generalizations of the ED paradigm.
+Binary decision tasks are selected since they can display our method most conveniently and clearly.
+Nevertheless, the proposed BAA mechanism is conceptually extensible to multi-value discrete decision tasks and applicable to a broader range of tasks involving discrete decisions, such as segmentation, classification, and object recognition.
+The main contributions of this work are summarized as follows:
+  • 1.
+    We identify a fundamental mismatch between continuous optimization and discrete inference in model training and propose the Binarization-Aware Adjuster (BAA) framework to bridge this gap.
+  • 2.
+    We develop a principled formulation based on the Distance Weight Function (DWF), enabling binarization-aware loss adjustment while retaining full differentiability.
+  • 3.
+    We apply the proposed framework to the edge detection task and implement extensive experiments on substantial benchmark datasets across multiple architectures.
+2 Previous Works
+A large number of learning-based systems perform discrete decision-making while being optimized using continuous-valued loss functions.
+This discrepancy has motivated studies on improving the alignment between continuous optimization during training and discrete evaluation at inference.
+In this work, edge detection (ED) is adopted as an illustrative validation scenario due to its strong dependency on binary decisions and its relevance to many structured prediction tasks.
+In the following, we briefly review related studies with ED as a representative example.
+2.1 Datasets
+Historically, early vision research commonly relied on general-purpose datasets.
+For instance, BSDS300/500 [14] have been used in contour detection, segmentation, and ED.
+Common datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have been widely adopted across segmentation, detection, and image classification tasks.
+With increasing task specialization, dataset selection became more focused, e.g., PASCAL VOC for object detection, NYUD for segmentation, and BSDS500 for contour-based analysis.
+For ED specifically, dedicated datasets were developed to provide more accurate edge annotations.
+BIPED/BIPED2 [18] offers edge annotations for 250 high-resolution campus images, BRIND [6] extended BSDS500 with reannotated edge labels, and UDED [19] provides carefully selected high-quality samples.
+Despite these efforts, annotation quality remains a major challenge.
+Human-labeled edge maps often suffer from inconsistency, ambiguity, and noise.
+Although robust training strategies and refined labeling protocols have been proposed [20, 21, 22], annotation noise still affects both generalization ability and fair benchmarking.
+2.2 Model Architectures
+Deep learning architectures have achieved broad success in computer vision [23], natural language processing [24], and decision systems [25].
+Major architectural paradigms, including convolutional neural networks (CNNs), transformer models, and diffusion-based methods, have demonstrated significant cross-domain applicability.
+Using ED as an example, the pioneering work HED [11] validated the effectiveness of multi-scale deep supervision using CNNs, outperforming traditional approaches [4, 5] and early feature-based learning methods [26, 27, 28, 29, 30].
+Subsequent works, such as RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further optimized fusion strategies and model structures.
+Transformer-based [7, 35] and diffusion-based [8, 9] designs have also been explored.
+More recently, the Extractor–Selector framework [10] demonstrated improvements through refined feature selection.
+Among these approaches, CNN-based models remain the most widely adopted due to their robustness and maturity.
+2.3 Loss Functions
+Loss function design is critical in model optimization.
+Binary cross-entropy (BCE) was initially the standard for decision tasks.
+However, BCE assigns equal importance to positive and negative samples, which is unsuitable for imbalanced data.
+To address this, Weighted BCE (WBCE) is widely adopted and still the most standard one currently.
+Other alternatives include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulation, and SWBCE [13] that models uncertainty in manual annotations.
+Although these methods improve sample weighting, they generally do not explicitly model the final inference rule, leaving the discretization step separate from the optimization process.
+Consequently, the fundamental continuous–discrete gap remains unresolved.
+This limitation motivates the development of our BAA framework.
+2.4 Evaluation Metrics
+[…]
+$F\beta=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
+where
+$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN}.$ (2)
+Here, TP denotes true positives, FP false positives, and FN false negatives, with $\beta$ typically set to 1. The most frequently used metrics include:
+  • 1.
+    ODS (Optimal Dataset Score): Best F1 using a single threshold for the entire dataset.
+  • 2.
+    OIS (Optimal Image Score): Mean of each image’s best F1 using image-specific thresholds.
+For ED, the commonly used metric computation was proposed in [26] with error tolerance to compensate for minor spatial deviations.
+Recent studies [22] recommend stricter and standardized tolerance settings for more reliable comparison.
+3 Methodology: General Theory
+This section presents the theoretical framework of our method.
+Here, the design for binary decision tasks is proposed for illustrative convenience and simplicity.
+Note that the discussions can be easily extended to multi-value discrete decision tasks.
+3.1 Motivation and Insight
+A key challenge in training binary decision models is the mismatch between training objectives and evaluation protocols.
+During training, models generate continuous-valued outputs, typically normalized to [0,1], and are optimized using pixel-wise loss functions such as WBCE.
+In contrast, evaluation metrics are computed on binarized maps, obtained by applying a fixed threshold to these outputs.
 Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process.
@@
 Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority.
-This misallocation of optimization effort can prevent the optimal parameter be learned, resulting in suboptimal learning and degraded final performance.
-For example, Table I reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8.
+This misallocation of optimization effort can prevent the optimal parameter from being learned, resulting in suboptimal learning and degraded final performance.
+For example, Table 1 reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8.
 Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss.
-Moreover, such confident correct predictions are often easy to optimize.
-Indeed, it is because they are easy to optimize that they become the correctly predicted ones with high confidence.
-Therefore, training naturally prioritizes them for their optimization ease, even though further improvement on them yields no benefit for the final decision.
+Moreover, such confident-correct predictions are often easy to optimize, which is the reason that they become the correctly predicted ones with high confidence.
+Therefore, training naturally prioritizes them for their optimization ease since they contribute more to loss, even though further improvement on them yields no benefit for the final decision.
 In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold.
 Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
-TABLE I: Optimal binarization thresholds under WBCE training: Evaluation uses 1-pixel error tolerance without NMS.
-Experimental settings follow the ones in Section IV-A.
+Table 1: Optimal binarization thresholds under WBCE training for ED models: Evaluation uses 1-pixel error tolerance without NMS.
+Experimental settings follow the ones in Section 4.1.
 Models are retrained on the pre-training set and evaluated on the validation set.
@@
 By explicitly incorporating binarization behavior into the training process, the BAA method ensures that the model allocates its learning capacity to the most decision-critical pixels.
-III-B Adjusted Loss Function Formulation
-Although designing the BAA is non-trivial, the idea of using an adjuster is conceptually simple.
+3.2 Adjusted Loss Function Formulation
+Although designing the BAA is non-trivial, the idea of using it is conceptually simple.
 Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
 We define an adjusted loss as a loss that incorporates a pixel-wise weighting factor:
-$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(P_{i},G_{i})+\delta)L(P_{i},G_{i})$ (1)
-where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single pixel ($L(P_{i},G_{i})$) and on the whole image ($L(Pred,Gt)$)., $Pred$ and $Gt$ represent the model prediction and the ground truth, respectively, $P_{i}$ and $G_{i}$ represent the model prediction and the ground truth of the $i$-th pixel, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the adjusted loss to stabilize training.
-The function $Adj$ is called an Adjuster, where $Adj(P_{i},G_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
-Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(P_{i},G_{i})$ is the binary mask $\hat{Adj}_{thr}(P_{i},G_{i})$, defined as:
-$\hat{Adj}_{thr}(P_{i},G_{i})=\left\{\begin{matrix}&0&If\ (P_{i}-thr)(G_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right.$ (2)
-where $thr$ is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
-  • If a pixel is correctly classified, then the loss on it is 0.
-  • If a pixel is incorrectly classified, then the loss on it is positive.
-This adjuster emphasizes only misclassified pixels.
+$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i})$ (3)
+where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single element such as a pixel ($L(Pred_{i},Gt_{i})$) and on the whole object such an images ($L(Pred,Gt)$)., PredPred and $Gt$ represent the model prediction and the ground truth, respectively, $Pred_{i}$ and $Gt_{i}$ represent the model prediction and the ground truth of the $i$-th element, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the loss to stabilize training.
+The function $Adj$ is called an Adjuster, where $Adj(Pred_{i},Gt_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
+Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(Pred_{i},Gt_{i})$ is the binary mask $\hat{Adj}_{thr}(Pred_{i},Gt_{i})$, defined as:
+$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&If\ (Pred_{i}-thr)(Gt_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right$. (4)
+where thrthr is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
+  • 1.
+    If an element is correctly classified, then the loss on it is 0.
+  • 2.
+    If an element is incorrectly classified, then the loss on it is positive.
+This adjuster emphasizes only misclassified elements.
 However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization.
 Therefore, a smoother, differentiable formulation is necessary.
-III-C Design of the Binarization-Aware Adjuster
-Let $thr$ denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
-  • $C1\ (Continuity)$: The adjuster $Adj$ must be continuous to allow gradient-based optimization.
-  • $C2\ (Non-negativity)$, $Adj(P_{i},G_{i})$ should be non-negative since it represents the weight of a pixel.
-Without loss generality, assume that $Adj(P_{i},G_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
-  • $C3\ (Uncertainty\ Emphasis)$: If a pixel is correctly classified, then the closer the prediction is to $thr$, the higher the weight $Adj(P_{i},G_{i})$ of it should be, reflecting the loss caused by the uncertainty.
-  • $C4\ (Confidence\ Saturation)$: If a pixel is correctly classified with high confidence (i.e., prediction far from $thr$), then $Adj(P_{i},G_{i})$ should decay to 0 quickly as $P_{i}$ goes away from $thr$, representing that further optimizing provide little benefits.
-Indeed, $Adj(P_{i},G_{i})$ should close to the maximal value when $P_{i}$ close to $thr$, remain large when $P_{i}$ is not far from $thr$, and decay to 0 after some pre-defined distance $thr\_dev$.
-  • $C5\ (Classification\ Sensitivity)$: If a pixel is misclassified, namely if it is an edge one but $P_{i}<thr$ or is a non-edge one but $P_{i}\geq thr$, its weight $Adj(P_{i},G_{i})$ should be maximal, representing correctly classification is of the first priority.
-These conditions ensure that misclassified pixels always receive full attention, and optimization of correct pixels is focused on the uncertain parts while conserving effort on highly confident correct parts.
+3.3 Design of the Binarization-Aware Adjuster
+Let thrthr denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
+  • 1.
+    C1(Continuity)C1\ (Continuity): The adjuster $Adj$ must be continuous to allow gradient-based optimization.
+  • 2.
+    C2(Nonnegativity)C2\ (Non-negativity), $Adj(Pred_{i},Gt_{i})$ should be non-negative since it represents the weight of an element.
+Without loss of generality, assume that $Adj(Pred_{i},Gt_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
+  • 3.
+    C3(UncertaintyEmphasis)C3\ (Uncertainty\ Emphasis): If an element is correctly classified, then the closer the prediction is to thrthr, the higher the weight $Adj(Pred_{i},Gt_{i})$ of it should be, reflecting the loss caused by the uncertainty.
+  • 4.
+    C4(ConfidenceSaturation)C4\ (Confidence\ Saturation): If a n element is correctly classified with high confidence (i.e., prediction far from thrthr), then $Adj(Pred_{i},Gt_{i})$ should decay to 0 quickly as $Pred_{i}$ goes away from thrthr, representing that further optimizing provide little benefits.
+Indeed, $Adj(Pred_{i},Gt_{i})$ should close to the maximal value when $Pred_{i}$ is close to thrthr, remain large when $Pred_{i}$ is not far from thrthr, and decay to 0 after some pre-defined distance thr_devthr\_dev.
+  • 5.
+    C5(ClassificationSensitivity)C5\ (Classification\ Sensitivity): If an element is misclassified, for example if it is an edge one but $Pred_{i}<thr$ or is a non-edge one but $Pred_{i}\geq thr$ in the ED task, its weight $Adj(Pred_{i},Gt_{i})$ should be maximal, representing correctly classification is of the first priority.
+These conditions ensure that misclassified elements always receive full attention, and optimization of correct elements is focused on the uncertain parts while conserving effort on highly confident correct parts.
 We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short.
@@
 Definition 1
-A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and $thr\_dev>0$ is a function $[0,1]^{2}\longrightarrow[0,1]$ satisfying the above conditions C1 to C5.
-Let $thr\_dev$ denote the threshold window from $C4$. To construct a BAA, we define that:
+A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued at [0,1] satisfying the above conditions C1 to C5.
+Let thr_devthr\_dev denote the threshold window from $C4$. To construct a BAA, we define:
 Definition 2
 A function $f_{thr\_dev}(x)$ is said to be a Distance Weight Function (DWF) regarding the threshold window $thr\_dev\in\mathbb{R}_{+}$, if it satisfies the following properties:
+  • 1.
     Continuity: $f_{thr\_dev}(x):[0,thr\_dev]\longrightarrow[0,1]$ is continuous.
+  • 2.
     Monotonically Decreasing: $f_{thr\_dev}(x_{1})\leq f_{thr\_dev}(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ with $x_{1}\geq x_{2}$.
 (or equivalently, $\frac{df_{thr\_dev}}{dx}(x)\leq 0$ for differentiable $f_{the\_dev}(x)$).
+  • 3.
     Concavity: $f(\lambda x_{1}+(1-\lambda)x_{2})\geq\lambda f(x_{1})+(1-\lambda)f(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ and $\lambda\in[0,1]$.
 (or equivalently, $\frac{d^{2}f_{thr\_dev}}{dx^{2}}(x)\leq 0$ for secondary differentiable $f_{the\_dev}(x)$).
+  • 4.
     Boundary Conditions: $f_{thr\_dev}(0)=1$, and $f_{thr\_dev}(thr\_dev)=0$.
 Then:
-$\bar{Adj}(P_{i},G_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|P_{i}-thr|)&0\leq|P_{i}-thr|\leq thr\_dev\\ &0\ &|P_{i}-thr|>thr\_dev\end{matrix}\right.$ (3)
-satisfies $C1-C4$.
+$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|Pred_{i}-thr|)&0\leq|Pred_{i}-thr|\leq thr\_dev\\ &0\ &|Pred_{i}-thr|>thr\_dev\end{matrix}\right$. (5)
+satisfies C1C4C1-C4.
 An example of the DWF $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ can be constructed as:
@@
 $MD_{thr}(x,y)=(x-thr)\times y+(thr-x)\times(1-y)$ (7)
-Hence, $MD_{thr}(P_{i},G_{i})$ measures both the correctness and confidence of the pixel $i$.
-For a wrongly predicted pixel $i$, namely whose prediction is on the different side of the $thr$ as its ground truth value, $MD_{thr}(P_{i},G_{i})\leq 0$, while for a correctly predicted pixel $i$, $MD_{thr}(P_{i},G_{i})$ is the $L^{1}$ distance of $P_{i}$ from the binarization threshold $thr$.
+Hence, $MD_{thr}(Pred_{i},Gt_{i})$ measures both the correctness and confidence of the element $i$.
+For a wrongly predicted element $i$, namely whose prediction is on the different side of the thrthr as its ground truth value, $MD_{thr}(Pred_{i},Gt_{i})\leq 0$, while for a correctly predicted element $i$, $MD_{thr}(Pred_{i},Gt_{i})$ is the $L^{1}$ distance of $Pred_{i}$ from the binarization threshold thrthr.
 Now, a BAA can be defined as:
-$BAA_{thr,thr\_dev}(P_{i},G_{i})=\hat{f}_{thr\_dev}(MD_{thr}(P_{i},G_{i}))$ (6)
+$BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=\hat{f}_{thr\_dev}(MD_{thr}(Pred_{i},Gt_{i}))$ (8)
 where $\hat{f}_{thr\_dev}$ is an extension of a selected DWF $f_{thr\_dev}(x)$, defined as:
 $\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<0\\ &f_{thr\_dev}(x)&0\leq x\leq thr\_dev\\ &0&x>thr\_dev\end{matrix}\right$. (9)
-III-D Theoretical Properties of the BAA
+3.4 Theoretical Properties of the BAA
 To better understand the proposed adjuster’s behavior, we analyze its mathematical properties.
@@
 Then $\hat{f}_{b,thr\_dev}(x)$ serves as a smooth approximation to a hard binarization function.
-Specifically, it generalizes the discrete threshold-based binary weighting function $B_{thr\_dev}(x)$ on threshold $thr\_dev$, defined as:
+Specifically, it generalizes the discrete threshold-based binary weighting function $B_{thr\_dev}(x)$ on threshold thr_devthr\_dev, defined as:
 $B_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<thr\_dev\\ &0\ &x\geq thr\_dev\end{matrix}\right$. (10)
@@
 $\lim_{b\rightarrow+\infty}\hat{f}_{b,thr\_dev}(x)=B_{thr\_dev}(x)$ (11)
-Therefore, the extended function $\hat{f}_{b,thr\_dev}$ is a continuous approximation for threshold logic $B_{the\_dev}(x)$.
-Accordingly, the map $BAA_{thr,thr\_dev}(P_{i},G_{i})=BAA_{thr,b,thr\_dev}(P_{i},G_{i})$, defined by Equation 6, converges to a discrete decision-based adjuster $\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$, defined as follow: {strip}
-$\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})=\left\{\begin{matrix}&0&(P_{i}-thr)(G_{i}-thr)\geq 0\ \&\ |P_{i}-thr|\geq thr\_dev\\ &1&Otherwise\end{matrix}\right.$ (10)
+Therefore, $\hat{f}_{b,thr\_dev}$ is a continuous approximation for threshold logic.
+Accordingly, the $BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=BAA_{thr,b,thr\_dev}(Pred_{i},Gt_{i})$, defined by Equation 8, converges to a discrete decision-based adjuster $\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})$, defined as:
+$\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&(Pred_{i}-thr)(Gt_{i}-thr)\geq 0\ \&\ |Pred_{i}-thr|\geq thr\_dev\\ &1&Otherwise\end{matrix}\right$. (12)
 i.e.,
-$\small\lim_{b\rightarrow+\infty}BAA_{thr,b,thr\_dev}(P_{i},G_{i})=\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$ (11)
+$\lim_{b\rightarrow+\infty}BAA_{thr,b,thr\_dev}(Pred_{i},Gt_{i})=\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})$ (13)
 This confirms that the BAA bridges the binarization logic with the continuous-loss gradient-based training.
@@
 It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided..
-These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation in ED.
-III-E Hyperparameter Selection and Training Protocol
-The proposed $BAA$ depends on three hyperparameters: $thr$: binarization threshold, $thr\_dev$: threshold window beyond which correct predictions are ignored in optimization, and $b$ (when employing $f_{b,thr\_dev}(x)$ from Equation 4 as the DWF): DWF decay rate,
-The DWF decay rate $b$ occurs specifically in the DWF defined in Equation 4, controlling the concavity of $f_{b,thr\_dev}(x)$, determining how sharply the adjusted weight decays as predictions diverge from the decision boundary $thr$.
+These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation.
+3.5 Hyperparameter Selection and Training Protocol
+The proposed $BAA$ depends on three hyperparameters: thrthr: binarization threshold, thr_devthr\_dev: threshold window beyond which correct predictions are ignored in optimization, and $b$ (when employing $f_{b,thr\_dev}(x)$ from Equation 6 as the DWF): DWF decay rate,
+The DWF decay rate $b$ occurs specifically in the DWF defined in Equation 6, controlling the concavity of $f_{b,thr\_dev}(x)$, determining how sharply the adjusted weight decays as predictions diverge from the decision boundary thrthr.
 Its effect is illustrated in Figure 1, and we set $b=16$ as the default.
 Figure 1: The different choices of $b$ in the DWF $f_{b,thr\_dev}(x)$ presented in Equation 6
-$thr\_dev$ defines the margin around $thr$ within which a correct prediction is considered uncertain.
-A correct prediction with $|P_{i}-thr|>thr\_dev$ is considered as confidently correct and receives zero weight.
-We empirically choose $thr\_dev=0.2$.
-$thr$ is the most critical hyperparameter in the BAA. Experiments (Table I) show that the optimal value for ED models typically lies between 0.6 and 0.8, so we use $thr=0.7$ as the default.
+thr_devthr\_dev defines the margin around thrthr within which a correct prediction is considered uncertain.
+A correct prediction with $|Pred_{i}-thr|>thr\_dev$ is considered as confidently correct and receives zero weight.
+We empirically choose thr_dev=0.2thr\_dev=0.2 for the ED task.
+thrthr is the most critical hyperparameter in the BAA. For ED models, experiments (Table 1) show that the optimal value typically lies between 0.6 and 0.8, so we use $thr=0.7$ as the default.
 However, given its importance, it is worth providing a calibration method for it.
-Specifically, we suggest a self-adaptive $thr$ chosen by pre-training: pre-train the model on a sub-dataset without applying BAA and evaluate the pre-trained model on a validation set to select a suitable $thr$ for BAA.
-Therefore, the final training protocol of an ED model $M$ under a selected baseline loss function $L$ on a dataset $D$, with the presented BAA method can be formulated as:
-  • Step 1: Split $D$ into pre-train set $PT$, validation set $V$, and test set $T$.
-  • Step 2: Train $M$ on $PT$ using baseline loss $L$, and evaluate on $V$ to estimate $thr$.
-  • Step 3: Retrain $M$ via $L_{BAA,\delta}$, on $PT\cup V$, and evaluate finally on $T$.
+Specifically, we suggest a self-adaptive thrthr chosen by pre-training: pre-train the model on a sub-dataset without applying BAA and evaluate the pre-trained model on a validation set to select a suitable thrthr for BAA.
+Therefore, the final training protocol of a model $M$ under a selected baseline loss function $L$ on a dataset $D$, with the presented BAA method, can be formulated as:
+  • 1.
+    Step 1: Split $D$ into pre-train set $P$, validation set $V$, and test set $T$.
+  • 2.
+    Step 2: Train $M$ on $P$ using baseline loss $L$, and evaluate on $V$ to estimate thrthr.
+  • 3.
+    Step 3: Retrain $M$ via $L_{BAA,\delta}$, on $P\cup V$, and evaluate finally on $T$.
 In experiments, $\delta$ is simply set to be 1.
-IV Experiment
-This section presents experiments designed to evaluate the proposed method.
-IV-A Experimental Setting
-To assess the effectiveness of the proposed binarization adjustment mechanism, we adopt the adjuster designed by Equation 4, 5, 6, and 7 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
-IV-A1 Baseline and Benchmark
-The proposed method is evaluated across a range of ED models, including:
-  • HED (2015) [11].
-  • BDCN (2022) [32].
-  • Dexi (2023) [34].
-  • EES3-Double (2025) [10], the EES framework, integrating the above three models with the double-size selector.
-Experiments are conducted on four benchmark datasets:
-  • BRIND [6].
-  • BSDS500 [14].
-  • UDED [22].
-  • BIPED2 [21].
+4 Experiment: Application to Edge Detection
+To validate the effectiveness of the proposed BAA framework, we apply it to the ED task, which is selected since it is one of the most fundamental tasks in computer vision, supporting substantial high-level tasks, and many tasks can be reformulated to it.
+4.1 Experimental Setting
+We adopt the adjuster designed by Equation 6, 7, 8, and 9 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
+4.1.1 Baseline and Benchmark
+The proposed method is evaluated across a range of ED models, including: HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], the EES framework integrating the above three models with the double-size selector.
+Experiments are conducted on four benchmark datasets: BRIND [6], BSDS500 [14], UDED [19], and BIPED2 [18].
 All models are retrained from scratch on each dataset to ensure fairness.
@@
 Each model is trained under the following three configurations of loss function for comparison:
+  • 1.
     Standard WBCE (baseline).
+  • 2.
     WBCE with a fixed-threshold BAA using $thr=0.7$ (denoted as BAA-0.7)
-  • WBCE with a self-adaptive adjuster where $thr$ is automatically selected following the method in Section III-E (denoted as BAA-SA).
-Unless otherwise specified, the hyperparameters for the adjusted loss defined by the BAA are fixed at $thr\_dev=0.2$, $b=16$, and $\delta=1$.
-Evaluation protocols follow the standard ED metrics (ODS and OIS) introduced in [26], under the strictest 1-pixel error tolerance without post-processing techniques such as non-maximal-suppression (NMS), suggested in [25].
+  • 3.
+    WBCE with a self-adaptive adjuster where thrthr is automatically selected following the method in Section 3.5 (denoted as BAA-SA).
+Unless otherwise specified, the hyperparameters for the adjusted loss defined by the BAA are fixed at thr_dev=0.2thr\_dev=0.2, $b=16$, and $\delta=1$.
+Evaluation protocols follow the standard ED metrics (ODS and OIS) introduced in [26], under the strictest 1-pixel error tolerance without post-processing techniques such as non-maximal suppression (NMS), suggested in [22].
 The evaluations under the traditional relaxed error tolerance (4.3 to 11.1 pixels) with NMS are provided in the appendix.
-IV-A2 Dataset Partitions and Augmentation
+4.1.2 Dataset Partitions and Augmentation
 Dataset preprocessing and postprocessing procedures follow those introduced in [10]. Details are stated as follows.
-The dataset partitions are:
-  • BRIND and BSDS500: 400 training images, 100 testing images.
-  • UDED: 20 training images, 7 testing images (excluding two images with resolution below $320\times 320$).
-  • BIPED2: 200 training images, 50 testing images.
-For training under the self-adaptive adjuster setting in section III-E, each training set is further divided into pre-training and validation subsets:
-  • BRIND and BSDS500: 300 pre-training, 100 validation images.
-  • UDED: 15 pre-training, 5 validation images.
-  • BIPED2: 150 pre-training, 50 validation images.
-Data augmentations are applied, including:
-  • Halving the images until both height and width are strictly below 640 pixels.
-  • Rotations at four angles, $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, with horizontal flips.
-  • Addition of noiseless data.
-IV-A3 Training Note
+The dataset partitions follow: BRIND and BSDS500: 400 training images and 100 testing images, UDED: 20 training images and 7 testing images (excluding two images with resolution below $320\times 320$), and BIPED2: 200 training images and 50 testing images.
+For training under the self-adaptive adjuster setting in section 3.5, each training set is further divided into pre-training and validation subsets, following: BRIND and BSDS500: 300 pre-training and 100 validation images, UDED: 15 pre-training and 5 validation images, and BIPED2: 150 pre-training and 50 validation images.
+Data augmentations are applied, including: Halving the images until both height and width are strictly below 640 pixels, rotations at four angles, $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, with horizontal flips, as well as addition of noiseless data.
+4.1.3 Training Note
 During training, input images are randomly cropped to $320\times 320$ and refreshed every 5 epochs.
@@
 Training runs for 200 epochs on UDED and 50 epochs on the remaining datasets.
-IV-A4 Inference Method
+4.1.4 Inference Method
 In the inference phase, each image is split into overlapping $320\times 320$ patches with a stride of 304 pixels (i.e., 16 pixels overlap).
 The model predicts edge maps on each patch independently, and the outputs are aggregated to form the final edge map.
-Further implementation details can be found in the codes or in [10].
-IV-B Experimental Result
-Tables II and III report the ED performance of all methods on BRIND, BSDS500, UDED, and BIPED2 datasets under the strict 1-pixel error tolerance setting, without applying NMS.
+Further implementation details can be found in the code or in [10].
+4.2 Experimental Results
+Tables 2 and 3 report the ED performance of all methods on BRIND, BSDS500, UDED, and BIPED2 datasets under the strict 1-pixel error tolerance setting, without applying NMS.
 The evaluation includes ODS and OIS scores for each configuration.
-TABLE II: Results on BRIND and BSDS500 under 1-pixel error tolerance without NMS: For BRIND, all annotations per image are merged into a single binary ground truth map, where a pixel is considered an edge one if it is marked as such by any annotator.
+Table 2: Results on BRIND and BSDS500 under 1-pixel error tolerance without NMS: For BRIND, all annotations per image are merged into a single binary ground truth map, where a pixel is considered an edge one if it is marked as such by any annotator.
 For BSDS500, only label 1 is used as the ground truth.
@@
 The best scores for each model are highlighted in bold.
-Rows labeled A, A-BAA-0.7, and A-BAA-SA represent the average performance across the four models, while Improvement rows indicate the relative improvements over the baseline.
+Rows labeled A-No-BAA, A-BAA-0.7, and A-BAA-SA represent the average performance across the four models corresponding to each setting, respectively, while Improvement rows indicate the relative improvements over the baseline.
 BRIND BSDS500
@@
 EES3-Double-BAA-SA 0.690 0.697 0.497 0.507
-A 0.669 0.677 0.457 0.470
+A-No-BAA 0.669 0.677 0.457 0.470
 A-BAA-0.7 0.674 0.682 0.464 0.475
@@
 […]
-TABLE III: Results on UDED and BIPED2 under 1-pixel error tolerance without NMS: Notations follow the ones in TableII.
+Table 3: Results on UDED and BIPED2 under 1-pixel error tolerance without NMS: Notations follow the ones in Table2.
 UDED BIPED2
@@
 EES3-Double-BAA-SA 0.758 0.791 0.675 0.680
-A 0.712 0.746 0.645 0.648
+A-No-BAA 0.712 0.746 0.645 0.648
 A-BAA-0.7 0.731 0.761 0.644 0.649
@@
 The method is robust across different backbone architectures and datasets, offering consistent improvements with minimal modifications to the existing training pipeline.
-IV-C Ablation Study
-Figure 2 illustrates how ODS and OIS scores vary with different values of $thr\_dev$ (from 0.1 to 0.7) and decay rates $b$ (8, 16, 32) in the DWF (Equation 4).
+4.3 Ablation Study
+Figure 2 illustrates how ODS and OIS scores vary with different values of thr_devthr\_dev (from 0.1 to 0.7) and decay rates $b$ (8, 16, 32) in the DWF (Equation 6).
 All experiments are conducted on the BRIND dataset using the EES3-Double model trained with the BAA-WBCE loss.
-The threshold $thr$ is fixed at 0.7, and $\delta$ at 1.
-This ablation study shows that across all tested configurations, $thr\_dev=0.2$ consistently yields the best performance.
+The threshold thrthr is fixed at 0.7, and $\delta$ at 1.
+This ablation study shows that across all tested configurations, thr_dev=0.2thr\_dev=0.2 consistently yields the best performance.
 Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
-Figure 2:  ODS and OIS across $thr\_dev$: Figures show how ODS and OIS vary with $thr\_dev$ values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32).
+Figure 2: ODS and OIS across thr_devthr\_dev: Figures show how ODS and OIS vary with thr_devthr\_dev values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32).
 Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss.
-Threshold $thr$ is fixed at 0.7 and $\delta$ at 1.
-V Conclusion
-This paper tackles a fundamental yet underexplored issue in image edge detection (ED): the misalignment between the continuous-valued outputs optimized during training and the binary predictions required at inference and evaluation.
-This discrepancy, often overlooked due to the discontinuity of binarization, undermines the alignment between training objectives and final task performance.
+Threshold thrthr is fixed at 0.7 and $\delta$ at 1.
+5 Discussion
+Although the BAA is initially designed for binary decision tasks, its formulation can be naturally extended to general multi-valued discrete inference.
+A straightforward strategy is to apply the adjuster to each decision boundary separately.
+That is, whenever a prediction approaches a discrete decision threshold, regardless of the number of categories, a higher optimization weight is assigned.
+The underlying theoretical principle remains consistent with the binary case established in Section 3, as each boundary can be locally treated as a binarization problem.
+While Section 4 presents experimental validation using ED as an example task, the proposed BAA is fundamentally applicable to a broader range of discrete prediction tasks.
+For instance, experimental results on the BSDS500 dataset, which is also used for contour detection and segmentation, implicitly demonstrate its broader potential.
+However, due to the substantial computational cost required for large-scale experiments [^3] The experiments on ED alone took approximately three months with GPUs.
+and the fact that this work primarily aims to establish the theoretical framework rather than exhaustive empirical study, we leave extensive cross-task evaluation for future research.
+Nevertheless, given that many structured prediction tasks in computer vision, such as segmentation, object detection, and boundary extraction, can be reformulated in terms of ED, we anticipate that the proposed method holds significant potential beyond.
+A limitation of the current self-adjusted BAA framework is that model training must be performed twice to determine the binarization threshold thrthr, incurring additional computational overhead during optimization, though it does not affect inference efficiency.
+Furthermore, the optimal decision threshold may vary across dataset partitions or benchmarks, requiring recalibration when transferring the model.
+This issue, however, is fundamentally induced by dataset-dependent evaluation settings rather than by the BAA framework itself.
+The self-adaptive threshold selection still provides a practical mechanism for choosing a reasonable threshold.
+Another drawback is that the self-estimated threshold may not precisely match the true optimal inference threshold.
+For example, in the ED task, the ideal threshold based on the ODS metric is obtained by test data, while the pre-trained threshold might be derived from it.
+Since the optimal threshold is inherently impossible to determine during training, the self-adaptive estimation should be regarded as a practically effective surrogate rather than an exact optimal value.
+Nevertheless, such a surrogate can also provide a method to estimate the real optimal decision threshold, since the real one is not practically accessible.
+In summary, although additional calibration is required and further task-specific validation remains open, the theoretical framework provides a principled foundation for improving optimization–inference alignment across a wide spectrum of discrete decision problems.
+6 Conclusion
+This paper tackles a fundamental yet underexplored issue occurring in discrete decision machine learning tasks: the misalignment between the continuous-valued outputs optimized during training and the discrete predictions required at inference and evaluation.
+This discrepancy, often overlooked due to the discontinuity of discretization, undermines the alignment between training objectives and final task performance.
 To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process.
-Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each pixel’s contribution to the loss based on its prediction correctness and confidence.
+Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each element’s contribution to the loss based on its prediction correctness and confidence.
 It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision.
@@
 In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior.
-Extensive experiments across diverse models and datasets validate the effectiveness and robustness of the proposed method.
-Beyond ED, our BAA approach can also provide a theoretical method for bridging the gap between continuous optimization and discrete evaluation, offering a promising direction for broader applications in structured prediction tasks such as segmentation and classification.
+Our method is applied to the edge detection (ED) task, and extensive experiments across representative models and datasets validate its effectiveness and robustness.
+In summary, the BAA introduces a theoretically grounded mechanism for bridging continuous optimization and discrete evaluation, with application to the ED task, demonstrating the effectiveness experimentally.
+And beyond ED task and binarization decision tasks, the principle can inform related discrete structured-prediction tasks, offering a promising direction for broader applications.

### v3

@@
-Binarization-Aware Adjuster: A Theoretical Framework for Bridging Continuous Optimization and Discrete Inference with Application to Edge Detection
+Binarization-Aware Adjuster for Discrete Decision Learning with an Application to Edge Detection
 Abstract
-In machine learning, discrete decision-making tasks exhibit a fundamental inconsistency between training and inference: models are optimized using continuous-valued outputs, yet evaluated through discrete predictions.
-This discrepancy arises from the non-differentiability of discretization operations, weakening the alignment between optimization objectives and practical decision outcomes.
-To address this, we present a theoretical framework for constructing a Binarization-Aware Adjuster (BAA) that integrates binarization behavior directly into gradient-based learning.
-Central to the approach is a Distance Weight Function (DWF) that dynamically modulates pixel-wise loss contributions based on prediction correctness and proximity to the decision boundary, thereby emphasizing decision-critical regions while de-emphasizing confidently correct samples.
-Furthermore, a self-adaptive threshold estimation procedure is introduced to better match optimization dynamics with inference conditions.
-As one of its applications, we implement experiments on the edge detection (ED) task, which also demonstrate the effectiveness of the proposed method experimentally.
-Beyond binary decision tasks and ED, the proposed framework provides a general strategy for aligning continuous optimization with discrete evaluation and can be extended to multi-valued decision processes in broader structured prediction problems.
+Discrete decision tasks in machine learning exhibit a fundamental misalignment between training and inference: models are optimized with continuous-valued outputs but evaluated using discrete predictions.
+This misalignment arises from the discontinuity of discretization operations, which prevents decision behavior from being directly incorporated into gradient-based optimization.
+To address this issue, we propose a theoretically grounded framework termed the Binarization-Aware Adjuster (BAA), which embeds binarization characteristics into continuous optimization.
+The framework is built upon the Distance Weight Function (DWF), which modulates loss contributions according to prediction correctness and proximity to the decision threshold, thereby aligning optimization emphasis with decision-critical regions while remaining compatible with standard learning pipelines.
+We apply the proposed BAA framework to the edge detection (ED) task, a representative binary decision problem.
+Experimental results on representative models and datasets show that incorporating BAA into optimization leads to consistent performance improvements, supporting its effectiveness.
+Overall, this work establishes a principled approach for aligning continuous optimization with discrete decision behavior, with its effectiveness demonstrated in a concrete application setting.
 keywords:
-Adjusted Loss Function, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
+Discrete Decision Learning, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
 […]
@@
 1 Introduction
-Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately require discrete decisions (e.g., binary or categorical outputs).
-However, learning methods, such as deep neural networks, constrained by the need for differentiable computation, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks [6, 7, 8, 9, 10].
-As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a fundamental gap arises between continuous-valued optimization during training and the discrete inference deployed at test time.
-This gap may lead to suboptimal performance, as optimization does not always align with decision-critical behavior.
-This mismatch is particularly evident in pixel-level binary prediction tasks.
+Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately involve discrete decision processes (e.g., binary or categorical outputs).
+However, modern learning methods, including deep neural networks, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks due to the requirement of differentiable computation [6, 7, 8, 9, 10].
+As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a practical inconsistency arises between continuous-valued optimization during training and discrete inference at test time.
+This inconsistency may lead to suboptimal learning behavior, as the optimization objective does not explicitly reflect decision-critical outcomes.
+This issue is particularly evident in pixel-level binary prediction tasks.
 For example, edge detection (ED) aims to generate binary edge maps but is typically trained using continuous-valued loss functions such as binary cross-entropy (BCE).
-Since binarization is non-differentiable, it is excluded from the optimization loop and applied only post hoc during evaluation.
-As a result, model updates tend to prioritize pixels that are easy to optimize but have limited influence on the binary output, whereas pixels near the decision boundary, which are more critical for final accuracy, receive insufficient emphasis due to their lower weights and higher optimization difficulty.
-Several studies have attempted to mitigate the inconsistency between decision behavior and training objectives, typically by changing losses that rebalance decision classes, such as weighted BCE [11] or alternative ones [12, 13].
-However, to the best of our knowledge, few existing methods explicitly integrate binarization (or discretization) behavior into gradient-based optimization, leaving the fundamental optimization–inference misalignment unresolved.
-To address this issue, we propose a theoretically grounded framework named Binarization-Aware Adjuster (BAA), which explicitly embeds binarization behavior into the optimization process.
-The core idea is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which dynamically reweights pixel-level loss according to prediction correctness and proximity to the decision threshold.
-This enables the model to allocate greater learning capacity to decision-critical regions while down-weighting confidently correct predictions that exert limited influence on binary outputs.
-Additionally, a self-adaptive threshold estimation strategy is introduced to align the adjuster with inference behavior.
-Since the BAA is continuous and differentiable, it can be seamlessly incorporated into existing gradient-based optimization pipelines, allowing discrete inference behavior to influence learning directly.
-The proposed framework is designed for binary decision tasks and is applied to the ED task to demonstrate its effectiveness.
-ED is selected as a representative implementation scenario due to its foundational role in low-level vision and its strong dependence on precise decision boundaries.
-Furthermore, many structured prediction tasks can be viewed as extensions or generalizations of the ED paradigm.
-Binary decision tasks are selected since they can display our method most conveniently and clearly.
-Nevertheless, the proposed BAA mechanism is conceptually extensible to multi-value discrete decision tasks and applicable to a broader range of tasks involving discrete decisions, such as segmentation, classification, and object recognition.
+Since binarization is non-differentiable, it is excluded from the optimization loop and applied only during inference and evaluation.
+Consequently, model updates tend to emphasize predictions that are easy to optimize but have limited impact on the final binary output, while predictions near the decision boundary, which are more critical for classification accuracy, often receive insufficient attention due to their lower contribution to the loss and higher optimization difficulty.
+Several studies have attempted to alleviate the discrepancy between training objectives and decision behavior, typically by modifying loss functions to rebalance decision classes, such as weighted BCE (WBCE) [11], or by adopting alternative loss formulations [12, 13].
+While these methods improve model performance, they do not explicitly model the influence of binarization on the optimization process.
+As a result, the underlying misalignment between continuous optimization and discrete inference remains largely unaddressed.
+To address this issue, we propose a theoretically motivated framework termed the Binarization-Aware Adjuster (BAA), which incorporates binarization-related decision characteristics into gradient-based optimization in a principled and continuous manner.
+The core idea of the BAA is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which reweights pixel-level loss contributions according to prediction correctness and proximity to the decision threshold.
+This formulation is designed to emphasize decision-critical predictions near the binarization boundary while reducing the influence of confidently correct predictions whose further optimization has limited effect on the final discrete outcome.
+Since the BAA is continuous, it can be directly integrated into standard training pipelines without altering model architectures or inference procedures.
+The proposed framework focuses on binary decision learning and is evaluated on the ED task as a representative application.
+ED is chosen due to its fundamental role in low-level vision and its strong dependence on precise binary decisions at the pixel level, making it well-suited for analyzing the interaction between continuous optimization and discrete inference.
+While the present study concentrates on ED, the formulation of the BAA is general in nature and provides theoretical insight into how decision boundaries can be more effectively reflected in continuous optimization objectives.
 The main contributions of this work are summarized as follows:
   • 1.
-    We identify a fundamental mismatch between continuous optimization and discrete inference in model training and propose the Binarization-Aware Adjuster (BAA) framework to bridge this gap.
+    We analyze a practical inconsistency between continuous-valued optimization and discrete inference in gradient-based learning and formalize its impact on binary prediction tasks.
   • 2.
-    We develop a principled formulation based on the Distance Weight Function (DWF), enabling binarization-aware loss adjustment while retaining full differentiability.
+    We propose the Binarization-Aware Adjuster (BAA), a principled and continuous loss adjustment framework based on the Distance Weight Function (DWF), to incorporate binarization behavior into the optimization process.
   • 3.
-    We apply the proposed framework to the edge detection task and implement extensive experiments on substantial benchmark datasets across multiple architectures.
+    We validate the proposed framework on the edge detection (ED) task through extensive experiments on multiple benchmark datasets and architectures.
 2 Previous Works
-A large number of learning-based systems perform discrete decision-making while being optimized using continuous-valued loss functions.
-This discrepancy has motivated studies on improving the alignment between continuous optimization during training and discrete evaluation at inference.
-In this work, edge detection (ED) is adopted as an illustrative validation scenario due to its strong dependency on binary decisions and its relevance to many structured prediction tasks.
-In the following, we briefly review related studies with ED as a representative example.
+A large body of learning-based systems perform discrete decision-making at inference while being optimized using continuous-valued loss functions during training.
+This discrepancy has motivated extensive research on improving the alignment between continuous optimization objectives and discrete evaluation criteria.
+In this work, edge detection (ED) is adopted as a representative validation scenario due to its strong dependence on binary decisions and its close connection to a wide range of structured prediction problems.
+In the following, we review relevant studies using ED as an illustrative example, with a focus on datasets, model architectures, loss functions, and evaluation protocols.
 2.1 Datasets
-Historically, early vision research commonly relied on general-purpose datasets.
-For instance, BSDS300/500 [14] have been used in contour detection, segmentation, and ED.
-Common datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have been widely adopted across segmentation, detection, and image classification tasks.
-With increasing task specialization, dataset selection became more focused, e.g., PASCAL VOC for object detection, NYUD for segmentation, and BSDS500 for contour-based analysis.
-For ED specifically, dedicated datasets were developed to provide more accurate edge annotations.
-BIPED/BIPED2 [18] offers edge annotations for 250 high-resolution campus images, BRIND [6] extended BSDS500 with reannotated edge labels, and UDED [19] provides carefully selected high-quality samples.
-Despite these efforts, annotation quality remains a major challenge.
-Human-labeled edge maps often suffer from inconsistency, ambiguity, and noise.
-Although robust training strategies and refined labeling protocols have been proposed [20, 21, 22], annotation noise still affects both generalization ability and fair benchmarking.
+Early research in computer vision commonly relied on general-purpose datasets.
+For instance, BSDS300/500 [14] have been widely used in contour detection, segmentation, and ED.
+Similarly, datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have supported a variety of tasks including segmentation, detection, and image classification.
+As research topics became more specialized, dataset construction increasingly focused on task-specific characteristics, e.g., PASCAL VOC for object detection, NYUD for indoor segmentation, and BSDS500 for contour-based analysis.
+For ED in particular, several dedicated datasets have been developed to provide more accurate and task-oriented annotations.
+BIPED and BIPED2 [18] provide high-resolution edge annotations for campus scenes, BRIND [6] extends BSDS500 with refined edge labels, and UDED [19] offers carefully curated samples with improved image and annotation quality.
+Despite these efforts, annotation quality remains a persistent challenge.
+Human-labeled edge maps often exhibit ambiguity, inconsistency, and noise, especially near weak or perceptually subjective boundaries.
+Although recent studies have proposed robust training strategies and improved annotation protocols [20, 21, 22], annotation noise continues to affect both model generalization and the reliability of quantitative evaluation.
 2.2 Model Architectures
-Deep learning architectures have achieved broad success in computer vision [23], natural language processing [24], and decision systems [25].
-Major architectural paradigms, including convolutional neural networks (CNNs), transformer models, and diffusion-based methods, have demonstrated significant cross-domain applicability.
-Using ED as an example, the pioneering work HED [11] validated the effectiveness of multi-scale deep supervision using CNNs, outperforming traditional approaches [4, 5] and early feature-based learning methods [26, 27, 28, 29, 30].
-Subsequent works, such as RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further optimized fusion strategies and model structures.
-Transformer-based [7, 35] and diffusion-based [8, 9] designs have also been explored.
-More recently, the Extractor–Selector framework [10] demonstrated improvements through refined feature selection.
-Among these approaches, CNN-based models remain the most widely adopted due to their robustness and maturity.
+Deep learning architectures have achieved remarkable success across a wide range of domains, including computer vision [23], natural language processing [24], and decision-making systems [25].
+Representative architectural paradigms, such as convolutional neural networks (CNNs) and transformers, have demonstrated strong modeling capacity and broad applicability.
+In the context of ED, the seminal work HED [11] demonstrated the effectiveness of multi-scale deep supervision within CNN architectures, significantly outperforming traditional edge detectors [4, 5] and earlier feature-based learning approaches [26, 27, 28, 29, 30].
+Subsequent CNN-based methods, including RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further refined network structures and feature fusion strategies.
+Transformer-based [7, 35] and diffusion-based [8, 9] approaches have also been explored to capture long-range dependencies and generative priors.
+More recently, the Extractor–Selector framework [10] introduced a modular design to improve feature selection.
+Despite architectural diversity, CNN-based models remain the most widely adopted in ED due to their robustness, efficiency, and mature training paradigms.
 2.3 Loss Functions
-Loss function design is critical in model optimization.
-Binary cross-entropy (BCE) was initially the standard for decision tasks.
-However, BCE assigns equal importance to positive and negative samples, which is unsuitable for imbalanced data.
-To address this, Weighted BCE (WBCE) is widely adopted and still the most standard one currently.
-Other alternatives include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulation, and SWBCE [13] that models uncertainty in manual annotations.
-Although these methods improve sample weighting, they generally do not explicitly model the final inference rule, leaving the discretization step separate from the optimization process.
-Consequently, the fundamental continuous–discrete gap remains unresolved.
-This limitation motivates the development of our BAA framework.
+Loss function design plays a central role in model optimization, particularly for binary decision tasks.
+BCE was initially the standard choice; however, it assigns equal importance to positive and negative samples, which is suboptimal for the highly imbalanced data distributions commonly encountered, such as in ED.
+To address this issue, WBCE has become a widely adopted baseline and remains one of the most standard loss formulations in practice.
+Beyond WBCE, various alternative loss functions have been proposed to better reflect structural or ranking-based objectives.
+Examples include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulations, and SWBCE [13], which explicitly models uncertainty in human annotations.
+While these approaches improve class balance, order cues, or perceptual quality, they generally treat binarization as a post-optimization operation and do not explicitly account for the inference rule during gradient-based learning.
+Consequently, the continuous–discrete inconsistency persists, motivating the development of the proposed BAA framework.
 2.4 Evaluation Metrics
 […]
-$F\beta=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
+$F_{\beta}=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
 where
 $\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN}.$ (2)
-Here, TP denotes true positives, FP false positives, and FN false negatives, with $\beta$ typically set to 1. The most frequently used metrics include:
+Here, $TP$ denotes true positives, $FP$ false positives, and $FN$ false negatives, with $\beta$ typically set to 1. The most commonly reported metrics include:
   • 1.
-    ODS (Optimal Dataset Score): Best F1 using a single threshold for the entire dataset.
+    ODS (Optimal Dataset Scale): The best F1 score achieved using a single optimal threshold across the dataset.
   • 2.
-    OIS (Optimal Image Score): Mean of each image’s best F1 using image-specific thresholds.
-For ED, the commonly used metric computation was proposed in [26] with error tolerance to compensate for minor spatial deviations.
-Recent studies [22] recommend stricter and standardized tolerance settings for more reliable comparison.
+    OIS (Optimal Image Scale): The average of image-wise best F1 scores obtained using image-specific optimal thresholds.
+For ED, the evaluation protocol with spatial tolerance was originally introduced in [26] to account for localization deviations.
+More recent studies [22] advocate stricter and standardized tolerance settings to improve the reliability and comparability of quantitative results.
 3 Methodology: General Theory
-This section presents the theoretical framework of our method.
-Here, the design for binary decision tasks is proposed for illustrative convenience and simplicity.
-Note that the discussions can be easily extended to multi-value discrete decision tasks.
+This section presents the theoretical framework of the proposed method.
+For clarity, the formulation is developed for binary decision tasks, which serve as a representative and intuitive setting.
+The discussion naturally extends to multi-valued discrete decision problems without altering the core principles.
 3.1 Motivation and Insight
-A key challenge in training binary decision models is the mismatch between training objectives and evaluation protocols.
-During training, models generate continuous-valued outputs, typically normalized to [0,1], and are optimized using pixel-wise loss functions such as WBCE.
-In contrast, evaluation metrics are computed on binarized maps, obtained by applying a fixed threshold to these outputs.
-Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process.
-This disconnect creates a fundamental issue: the loss function does not explicitly align with the discrete classification decisions that ultimately determine evaluation performance.
-This gap introduces a critical inefficiency in optimization.
-Pixels whose predicted values lie on the correct side and far from the decision threshold, namely those correct pixels with high confidence, tend to dominate the gradient signal since they are easy to optimize.
-In contrast, pixels with predictions near the threshold, which are more ambiguous and critical to classification, contribute less to the gradients since they are more difficult to optimize.
-Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority.
-This misallocation of optimization effort can prevent the optimal parameter from being learned, resulting in suboptimal learning and degraded final performance.
-For example, Table 1 reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8.
-Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss.
-Moreover, such confident-correct predictions are often easy to optimize, which is the reason that they become the correctly predicted ones with high confidence.
-Therefore, training naturally prioritizes them for their optimization ease since they contribute more to loss, even though further improvement on them yields no benefit for the final decision.
-In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold.
-Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
+A fundamental challenge in training binary decision models lies in the inconsistency between training objectives and evaluation protocols.
+During training, model outputs are continuous-valued, typically normalized to $[0,1]$, and optimized using pixel-wise loss functions such as WBCE.
+In contrast, evaluation metrics are computed on binarized outputs obtained by applying a fixed threshold.
+Since binarization is a discontinuous operation, it cannot be directly incorporated into gradient-based optimization.
+As a result, the loss function does not explicitly reflect the discrete decisions that ultimately determine evaluation performance.
+This inconsistency leads to an inefficiency in optimization.
+Predictions that are already on the correct side of the decision threshold and far from it, namely, confidently correct elements, tend to dominate the gradient signal because they are easy to optimize and contribute substantially to the loss.
+In contrast, predictions near the decision boundary, which are more ambiguous and critical for correct classification, are harder to optimize and therefore contribute less to the gradients.
+Consequently, the optimizer is implicitly biased toward improving easy but decision-irrelevant predictions, while allocating insufficient learning capacity to ambiguous yet decision-critical ones.
+This imbalance can prevent the model from converging to parameters that are optimal with respect to the final binary decisions.
+As an illustrative example, Table 1 reports optimal binarization thresholds for several ED models trained with WBCE across multiple datasets.
+These thresholds typically lie between $0.6$ and $0.8$.
+Suppose the threshold is set to $0.7$.
+Reducing the prediction of a non-edge pixel from $0.50$ to $0.01$ can significantly decrease the loss, yet it has no effect on the final binary edge map.
+Such predictions are often optimized early in training and tend to dominate the optimization process due to their large quantity and easily reducible loss values.
+In contrast, increasing the prediction of an edge pixel from $0.69$ to $0.71$ directly alters the binary decision outcome, but is generally more difficult to optimize as a pixel with the predicted value close to the decision threshold.
+Moreover, these pixels typically appear near true edge regions, which are more critical for accurate discrimination, but are relatively sparse in number.
+As a result, their overall contribution to the loss function is limited, making such decision-critical predictions more likely to be under-emphasized during training.
 Table 1: Optimal binarization thresholds under WBCE training for ED models: Evaluation uses 1-pixel error tolerance without NMS.
@@
 EES3-Double 0.61 0.86 0.77 0.80
-To bridge the continuous optimization and the binarization inference, we propose a novel loss adjustment mechanism, realized by a BAA, that dynamically re-weights pixel-wise loss contributions based on both proximity to the decision threshold and classification correctness.
-By explicitly incorporating binarization behavior into the training process, the BAA method ensures that the model allocates its learning capacity to the most decision-critical pixels.
+To address this issue, we propose a loss adjustment mechanism realized by the Binarization-Aware Adjuster (BAA), which dynamically reweights pixel-wise loss contributions based on both classification correctness and proximity to the decision threshold.
+By explicitly incorporating binarization-related decision characteristics into the optimization process, the proposed approach encourages the model to allocate greater learning capacity to decision-critical elements.
 3.2 Adjusted Loss Function Formulation
-Although designing the BAA is non-trivial, the idea of using it is conceptually simple.
-Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
-We define an adjusted loss as a loss that incorporates a pixel-wise weighting factor:
-$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i})$ (3)
-where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single element such as a pixel ($L(Pred_{i},Gt_{i})$) and on the whole object such an images ($L(Pred,Gt)$)., PredPred and $Gt$ represent the model prediction and the ground truth, respectively, $Pred_{i}$ and $Gt_{i}$ represent the model prediction and the ground truth of the $i$-th element, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the loss to stabilize training.
-The function $Adj$ is called an Adjuster, where $Adj(Pred_{i},Gt_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
-Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(Pred_{i},Gt_{i})$ is the binary mask $\hat{Adj}_{thr}(Pred_{i},Gt_{i})$, defined as:
-$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&If\ (Pred_{i}-thr)(Gt_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right$. (4)
-where thrthr is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
+We begin by formulating the adjusted loss function, deferring the specific construction of the BAA to the subsequent subsection.
+The adjusted loss is defined as:
+$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i}),$ (3)
+where $i$ indexes pixels, $L$ denotes a baseline pixel-wise loss such as WBCE [^1] The same symbol $L$ is used to denote the loss on a single element $L(Pred_{i},Gt_{i})$ and on the entire prediction $L(Pred,Gt)$., PredPred and $Gt$ represent the model prediction and ground truth, respectively, and $\delta\geq 0$ is a constant added to stabilize training.
+The function $Adj$, referred to as an adjuster, assigns dynamic weights in $[0,1]$ according to each element’s contribution to the final binary decision.
+A straightforward but impractical choice of $Adj$ is the hard binary mask:
+$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}0&\text{if }(Pred_{i}-thr)(Gt_{i}-thr)\geq 0,\\ 1&\text{otherwise},\end{matrix}\right$. (4)
+where thrthr denotes the binarization threshold.
+This formulation ensures that correctly classified elements incur zero loss, while misclassified ones incur positive loss, directly reflecting the binary decision.
+However, the resulting loss is discontinuous and unsuitable for gradient-based optimization, motivating the need for a continuous approximation.
+3.3 Design of the Binarization-Aware Adjuster
+Let thrthr denote the binarization threshold. The adjuster $Adj$ is designed to satisfy the following conditions:
   • 1.
-    If an element is correctly classified, then the loss on it is 0.
+    C1 (Continuity): $Adj$ must be continuous to support gradient-based optimization.
   • 2.
-    If an element is incorrectly classified, then the loss on it is positive.
-This adjuster emphasizes only misclassified elements.
-However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization.
-Therefore, a smoother, differentiable formulation is necessary.
-3.3 Design of the Binarization-Aware Adjuster
-Let thrthr denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
-  • 1.
-    C1(Continuity)C1\ (Continuity): The adjuster $Adj$ must be continuous to allow gradient-based optimization.
-  • 2.
-    C2(Nonnegativity)C2\ (Non-negativity), $Adj(Pred_{i},Gt_{i})$ should be non-negative since it represents the weight of an element.
-Without loss of generality, assume that $Adj(Pred_{i},Gt_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
+    C2 (Non-negativity): $Adj(Pred_{i},Gt_{i})\in[0,1]$, as it represents a weighting factor.
   • 3.
-    C3(UncertaintyEmphasis)C3\ (Uncertainty\ Emphasis): If an element is correctly classified, then the closer the prediction is to thrthr, the higher the weight $Adj(Pred_{i},Gt_{i})$ of it should be, reflecting the loss caused by the uncertainty.
+    C3 (Uncertainty Emphasis): For correctly classified elements, the weight should decrease as the prediction goes away from thrthr.
   • 4.
-    C4(ConfidenceSaturation)C4\ (Confidence\ Saturation): If a n element is correctly classified with high confidence (i.e., prediction far from thrthr), then $Adj(Pred_{i},Gt_{i})$ should decay to 0 quickly as $Pred_{i}$ goes away from thrthr, representing that further optimizing provide little benefits.
-Indeed, $Adj(Pred_{i},Gt_{i})$ should close to the maximal value when $Pred_{i}$ is close to thrthr, remain large when $Pred_{i}$ is not far from thrthr, and decay to 0 after some pre-defined distance thr_devthr\_dev.
+    C4 (Confidence Saturation): For confidently correct elements far from thrthr, the weight should rapidly decay, become zero beyond a predefined distance thr_devthr\_dev.
   • 5.
-    C5(ClassificationSensitivity)C5\ (Classification\ Sensitivity): If an element is misclassified, for example if it is an edge one but $Pred_{i}<thr$ or is a non-edge one but $Pred_{i}\geq thr$ in the ED task, its weight $Adj(Pred_{i},Gt_{i})$ should be maximal, representing correctly classification is of the first priority.
-These conditions ensure that misclassified elements always receive full attention, and optimization of correct elements is focused on the uncertain parts while conserving effort on highly confident correct parts.
-We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short.
-Namely:
+    C5 (Classification Sensitivity): Misclassified elements should receive the maximal weight.
+An adjuster satisfying these properties is referred to as a Binarization-Aware Adjuster (BAA).
 Definition 1
-A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued at [0,1] satisfying the above conditions C1 to C5.
-Let thr_devthr\_dev denote the threshold window from $C4$. To construct a BAA, we define:
+A Binarization-Aware Adjuster (BAA) with respect to $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued in $[0,1]$ satisfying conditions C1–C5.
+To construct such an adjuster, we introduce the Distance Weight Function (DWF).
 Definition 2
-A function $f_{thr\_dev}(x)$ is said to be a Distance Weight Function (DWF) regarding the threshold window $thr\_dev\in\mathbb{R}_{+}$, if it satisfies the following properties:
-  • 1.
-    Continuity: $f_{thr\_dev}(x):[0,thr\_dev]\longrightarrow[0,1]$ is continuous.
-  • 2.
-    Monotonically Decreasing: $f_{thr\_dev}(x_{1})\leq f_{thr\_dev}(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ with $x_{1}\geq x_{2}$.
-(or equivalently, $\frac{df_{thr\_dev}}{dx}(x)\leq 0$ for differentiable $f_{the\_dev}(x)$).
-  • 3.
-    Concavity: $f(\lambda x_{1}+(1-\lambda)x_{2})\geq\lambda f(x_{1})+(1-\lambda)f(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ and $\lambda\in[0,1]$.
-(or equivalently, $\frac{d^{2}f_{thr\_dev}}{dx^{2}}(x)\leq 0$ for secondary differentiable $f_{the\_dev}(x)$).
-  • 4.
-    Boundary Conditions: $f_{thr\_dev}(0)=1$, and $f_{thr\_dev}(thr\_dev)=0$.
-Then:
-$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|Pred_{i}-thr|)&0\leq|Pred_{i}-thr|\leq thr\_dev\\ &0\ &|Pred_{i}-thr|>thr\_dev\end{matrix}\right$. (5)
-satisfies C1C4C1-C4.
-An example of the DWF $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ can be constructed as:
-$f_{b,thr\_dev}(x)=\frac{e^{b\times x}-e^{b\times thr\_dev}}{1-e^{b\times thr\_dev}}$ (6)
-where $b\geq 0$ is pre-defined, controlling the decay rate.
-To also meet $C5$, we define the masked distance as:
+A function $f_{thr\_dev}(x)$ is a Distance Weight Function (DWF) with threshold window $thr\_dev\in\mathbb{R}_{+}$ if it satisfies continuity, monotonic decrease, concavity, and boundary conditions: $f_{thr\_dev}(0)=1$ and $f_{thr\_dev}(thr\_dev)=0$.
+Using a DWF, an intermediate adjuster satisfying C1–C4 can be defined as:
+$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}f_{thr\_dev}(|Pred_{i}-thr|),&|Pred_{i}-thr|\leq thr\_dev,\\ 0,&|Pred_{i}-thr|>thr\_dev.\end{matrix}\right$. (5)
+A concrete example of DWF is:
+$f_{thr\_dev}(x):=f_{b,thr\_dev}(x):=\frac{e^{b\times x}-e^{b\times thr\_dev}}{1-e^{b\times thr\_dev}},$ (6)
+where $b\geq 0$ controls the decay rate.
+To additionally satisfy C5, we define the masked distance:
 Definition 3
@@
 For a wrongly predicted element $i$, namely whose prediction is on the different side of the thrthr as its ground truth value, $MD_{thr}(Pred_{i},Gt_{i})\leq 0$, while for a correctly predicted element $i$, $MD_{thr}(Pred_{i},Gt_{i})$ is the $L^{1}$ distance of $Pred_{i}$ from the binarization threshold thrthr.
-Now, a BAA can be defined as:
-$BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=\hat{f}_{thr\_dev}(MD_{thr}(Pred_{i},Gt_{i}))$ (8)
+The final BAA is then defined as:
+$BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=\hat{f}_{thr\_dev}(MD_{thr}(Pred_{i},Gt_{i})),$ (8)
 where $\hat{f}_{thr\_dev}$ is an extension of a selected DWF $f_{thr\_dev}(x)$, defined as:
-$\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<0\\ &f_{thr\_dev}(x)&0\leq x\leq thr\_dev\\ &0&x>thr\_dev\end{matrix}\right$. (9)
+$\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}1,&x<0,\\ f_{thr\_dev}(x),&0\leq x\leq thr\_dev,\\ 0,&x>thr\_dev.\end{matrix}\right$. (9)
 3.4 Theoretical Properties of the BAA
@@
 In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 6, generalizes the baseline loss $L$.
-It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided..
+It degenerates to the baseline loss (up to a scale factor) when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains continuity.
 These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation.
@@
 4 Experiment: Application to Edge Detection
-To validate the effectiveness of the proposed BAA framework, we apply it to the ED task, which is selected since it is one of the most fundamental tasks in computer vision, supporting substantial high-level tasks, and many tasks can be reformulated to it.
+To validate the effectiveness of the proposed BAA framework, we apply it to the ED task.
+ED is selected as it is a fundamental problem in computer vision, serving as a core component for many higher-level vision tasks and providing a clear setting for evaluating binary decision learning.
 4.1 Experimental Setting
-We adopt the adjuster designed by Equation 6, 7, 8, and 9 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
+We use weighted binary cross-entropy (WBCE) as the baseline loss function $L$, due to its widespread adoption in ED tasks, and adopt the adjuster defined in Equations 6, 7, 8, and 9 as the BAA.
 4.1.1 Baseline and Benchmark
-The proposed method is evaluated across a range of ED models, including: HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], the EES framework integrating the above three models with the double-size selector.
+The proposed method is evaluated on a representative set of ED models, including HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], where EES3-Double denotes the EES framework integrating the above three models with a double-size selector.
 Experiments are conducted on four benchmark datasets: BRIND [6], BSDS500 [14], UDED [19], and BIPED2 [18].
@@
 For consistency and simplicity, only the final output of each model is used for supervision, bypassing any need for hyperparameter tuning in intermediate layers.
-Each model is trained under the following three configurations of loss function for comparison:
+Each model is trained under the following three loss configurations for comparison:
   • 1.
@@
 4.1.2 Dataset Partitions and Augmentation
-Dataset preprocessing and postprocessing procedures follow those introduced in [10]. Details are stated as follows.
+Dataset preprocessing and postprocessing procedures follow those introduced in [10]. The details are summarized as follows.
 The dataset partitions follow: BRIND and BSDS500: 400 training images and 100 testing images, UDED: 20 training images and 7 testing images (excluding two images with resolution below $320\times 320$), and BIPED2: 200 training images and 50 testing images.
@@
 […]
-The results clearly demonstrate that incorporating the proposed BAA method, whether with the fixed threshold (BAA-0.7) or the self-adaptive strategy (BAA-SA), leads to improvements in performance across most models and datasets.
-Both variants of the BAA loss improve the baseline WBCE loss across all models and datasets.
-The self-adaptive version (BAA-SA) tends to achieve slightly better performance than the fixed threshold in most cases.
+The results indicate that incorporating the proposed BAA method, either with a fixed threshold (BAA-0.7) or with the self-adaptive setting (BAA-SA), generally leads to improved performance across most evaluated models and datasets.
+Both variants of the BAA loss outperform the baseline WBCE loss in the majority of evaluated settings.
+The self-adaptive variant (BAA-SA) achieves slightly higher scores than the fixed-threshold version in several cases.
 […]
-These results strongly validate the effectiveness of the proposed BAA method, especially the self-adaptive variant.
-The method is robust across different backbone architectures and datasets, offering consistent improvements with minimal modifications to the existing training pipeline.
+Overall, these results support the effectiveness of the proposed BAA method in the context of ED.
+The performance gains are consistent across different backbone architectures and datasets, while requiring no modification to model architectures or inference procedures.
 4.3 Ablation Study
@@
 This ablation study shows that across all tested configurations, thr_dev=0.2thr\_dev=0.2 consistently yields the best performance.
-Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
-Figure 2: ODS and OIS across thr_devthr\_dev: Figures show how ODS and OIS vary with thr_devthr\_dev values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32).
-Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss.
-Threshold thrthr is fixed at 0.7 and $\delta$ at 1.
+Moreover, the proposed BAA method is robust to a wide range of hyperparameter settings, suggesting that it is not overly sensitive to hyperparameter choices in the ED setting and that a unified configuration can be adopted in practice.
+Figure 2: ODS and OIS across thr_devthr\_dev: ODS and OIS values under different thr_devthr\_dev settings (0.1 to 0.7) and decay rates $b$ (8, 16, 32).
+Experiments are conducted on BRIND using the EES3-Double model with the adjusted WBCE loss.
+The threshold thrthr is fixed at 0.7 and $\delta$ at 1.
 5 Discussion
-Although the BAA is initially designed for binary decision tasks, its formulation can be naturally extended to general multi-valued discrete inference.
-A straightforward strategy is to apply the adjuster to each decision boundary separately.
-That is, whenever a prediction approaches a discrete decision threshold, regardless of the number of categories, a higher optimization weight is assigned.
-The underlying theoretical principle remains consistent with the binary case established in Section 3, as each boundary can be locally treated as a binarization problem.
-While Section 4 presents experimental validation using ED as an example task, the proposed BAA is fundamentally applicable to a broader range of discrete prediction tasks.
-For instance, experimental results on the BSDS500 dataset, which is also used for contour detection and segmentation, implicitly demonstrate its broader potential.
-However, due to the substantial computational cost required for large-scale experiments [^3] The experiments on ED alone took approximately three months with GPUs.
-and the fact that this work primarily aims to establish the theoretical framework rather than exhaustive empirical study, we leave extensive cross-task evaluation for future research.
-Nevertheless, given that many structured prediction tasks in computer vision, such as segmentation, object detection, and boundary extraction, can be reformulated in terms of ED, we anticipate that the proposed method holds significant potential beyond.
-A limitation of the current self-adjusted BAA framework is that model training must be performed twice to determine the binarization threshold thrthr, incurring additional computational overhead during optimization, though it does not affect inference efficiency.
-Furthermore, the optimal decision threshold may vary across dataset partitions or benchmarks, requiring recalibration when transferring the model.
-This issue, however, is fundamentally induced by dataset-dependent evaluation settings rather than by the BAA framework itself.
-The self-adaptive threshold selection still provides a practical mechanism for choosing a reasonable threshold.
-Another drawback is that the self-estimated threshold may not precisely match the true optimal inference threshold.
-For example, in the ED task, the ideal threshold based on the ODS metric is obtained by test data, while the pre-trained threshold might be derived from it.
-Since the optimal threshold is inherently impossible to determine during training, the self-adaptive estimation should be regarded as a practically effective surrogate rather than an exact optimal value.
-Nevertheless, such a surrogate can also provide a method to estimate the real optimal decision threshold, since the real one is not practically accessible.
-In summary, although additional calibration is required and further task-specific validation remains open, the theoretical framework provides a principled foundation for improving optimization–inference alignment across a wide spectrum of discrete decision problems.
+Although the BAA is formulated for binary decision tasks, its underlying principle is not restricted to binarization itself.
+Conceptually, the framework can be extended to multi-valued discrete inference by applying the adjuster around individual decision boundaries.
+In such cases, when a correct prediction goes away from any discrete boundary, a decreased optimization weight can be assigned accordingly.
+From a theoretical perspective, this extension can even be realized by the binary formulation established in Section 3, as each local decision boundary can be treated as a binary decision problem.
+In this work, experimental validation is conducted using ED as a representative task.
+This choice is motivated by the clear binary inference rule in ED and its suitability for analyzing optimization-inference misalignment.
+The primary goal of this study is to establish a theoretically grounded adjustment mechanism rather than to provide an exhaustive empirical evaluation across tasks.
+While ED also underlies several related vision problems, we deliberately limit our experiments to this setting to maintain a focused and controlled analysis.
+Further empirical investigations on other tasks are therefore left as future work.
+One practical consideration of the BAA formulation is that additional model training for the binarization threshold estimation introduces additional training cost, despite not affecting inference efficiency.
+However, this overhead is not inherent to the BAA framework itself, but rather stems from the fact that the optimal decision threshold is not accessible in practical tasks.
+From this perspective, the two-stage estimation process should also be regarded as a general-purpose strategy for approximating decision thresholds in discrete inference problems, rather than as a limitation specific to the BAA framework.
+Another limitation is that the estimated threshold may not coincide exactly with the true optimal inference threshold.
+For instance, in ED, the optimal threshold under the ODS metric is determined using test data, which is inherently inaccessible during training.
+As a result, the self-adaptive threshold should be interpreted as an effective surrogate rather than an exact optimum.
+Nevertheless, such a surrogate provides a principled and practically feasible approximation for guiding binarization-aware optimization.
+Overall, despite the need for threshold calibration and the focus on a single task, the proposed framework establishes a principled foundation for improving the alignment between continuous optimization and discrete inference in decision-based learning problems.
 6 Conclusion
-This paper tackles a fundamental yet underexplored issue occurring in discrete decision machine learning tasks: the misalignment between the continuous-valued outputs optimized during training and the discrete predictions required at inference and evaluation.
-This discrepancy, often overlooked due to the discontinuity of discretization, undermines the alignment between training objectives and final task performance.
-To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process.
-Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each element’s contribution to the loss based on its prediction correctness and confidence.
-It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision.
-The resulting adjusted loss function from the BAA generalizes traditional loss formulations and embeds binarization behavior directly into the optimization process.
-In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior.
-Our method is applied to the edge detection (ED) task, and extensive experiments across representative models and datasets validate its effectiveness and robustness.
-In summary, the BAA introduces a theoretically grounded mechanism for bridging continuous optimization and discrete evaluation, with application to the ED task, demonstrating the effectiveness experimentally.
-And beyond ED task and binarization decision tasks, the principle can inform related discrete structured-prediction tasks, offering a promising direction for broader applications.
+This paper addresses a fundamental issue in discrete decision learning: the misalignment between continuous-valued optimization during training and discrete predictions at inference, which arises from the discontinuity of discretization operations that are excluded from standard learning pipelines.
+To address this issue, we introduce a theoretically grounded binarization-aware adjustment (BAA) mechanism, guided by the Distance Weight Function (DWF).
+BAA incorporates binarization characteristics into continuous optimization by modulating loss contributions according to prediction correctness and proximity to the decision threshold, emphasizing decision-critical elements and downweighting those with limited influence.
+The proposed framework is applied to the edge detection (ED) task, where binary inference is misaligned with the standard model training.
+Experimental results on representative models and datasets show that adopting BAA into optimization consistently improves performance.
+Overall, the BAA provides a principled mechanism for bridging continuous optimization and discrete evaluation, offering a theoretically sound direction for improving decision-aligned learning in discrete inference tasks.